### v1

# A Covariant and Manifestly Projective Invariant Formulation of Thomas-Whitehead Gravity

A Covariant and Manifestly Projective Invariant Formulation of Thomas-Whitehead Gravity
[…]
${}^{a}$ Department of Physics and Astronomy
The University of Iowa, Iowa City, IA 52242, USA
${}^{b}$ Department of Mathematics
University of California, Santa Barbara, CA 93106-3080
Thomas-Whitehead (TW) gravity is a recently formulated projectively invariant extension of Einstein-Hilbert gravity. Projective geometry was used long ago by Thomas et. al. to succinctly package equivalent paths encoded by the geodesic equation. Projective invariance in gravity has further origins in string theory through a geometric action constructed from the method of coadjoint orbits using the Virasoro Algebra. A projectively invariant connection arises from this construction, a part of which is known as the diffeomorphism field. TW Gravity exploits projective Gauss-Bonnet terms in the action functional to endow the diffeomorphism field with dynamics, while allowing the theory to collapse to General Relativity in the limit that the diffeomorphism field vanishes and the connection becomes Levi-Civita. In the original formulation of TW gravity, the diffeomorphism field is projectively invariant but not tensorial and the connection is projectively invariant but not affine. In this paper we reformulate TW gravity in terms of projectively invariant tensor fields and a projectively invariant covariant derivative, derive field equations respecting these symmetries, and show that the field equations obtained are classically equivalent across formulations. This provides a ’Rosetta Stone’ between this newly constructed covariant and projective invariant formulation of TW gravity and the original formulation that was manifestly projectively invariant, but not covariant.
1 Introduction
Nearly a century ago [1, 2], Thomas used projective geometry to package the set of equivalent paths encoded by the geodesic equation into a projectively invariant framework. In these seminal works, the connection is extended to be projectively invariant via the addition of a projective direction to make projective invariance manifest. This extra projective dimension was subsequently utilized by Veblin and Hoffman [3] in an attempt at unification of gravity with electromagnetism $\grave{\text{a}}$ la Kaluza and Klein [4, 5]. Whitehead further generalized the work of Thomas by allowing certain components of the projective connection to be arbitrary [6].
In the past few decades Whitehead’s generalization has been revisited [7, 8, 9, 10, 11] and projective geometry has been investigated in regards to the equivalence principle and cosmology [12, 13, 14, 15] and Einstein metrics in general [16, 17]. In making projective invariance manifest, the aforementioned projective direction extends the ${\rm d}$-dimensional base manifold $\mathcal{M}$ to a ${\rm d}+1$-dimensional manifold $\mathcal{N}$ now known as the Thomas cone. Thomas-Whitehead (TW) gravity is a recently formulated model that extends Einstein-Hilbert gravity to the Thomas cone [18, 19, 20]. Additional dynamical fields are introduced to perform this extension. The resulting TW gravity is projectively invariant and reduces to Einstein-Hilbert gravity upon setting the additional dynamical fields to zero.
In TW gravity, the additional projective direction is integrated out. This leads back to a model on the ${\rm d}$-dimensional manifold $\mathcal{M}$, where the additional fields take the form of a rank-two object $\mathcal{D}_{ab}$ known as the diffeomorphism field and a projective connection $\Pi^{c}{}_{ab}$. The connection and the metric are independent in the spirit of the Palatini formalism [21, 22], allowing the field equations to collapse naturally to the Einstein equations when the diffeomorphism field vanishes and the connection is Levi-Civita. In this way projective geometry can influence Riemannian geometry by acting as sources in the energy-momentum tensor, providing an avenue for geometric explanations of dark energy, dark matter, inflation, and other physical phenomenon. Indeed, the diffeomorphism field has recently been shown to be a possible source of dark energy [19] and inflation [23]. All stemming from the principle of projective invariance.
The roots of TW gravity go deeper still, to string theory and the Virasoro algebra. The notion of the diffeomorphism field as a dynamical field arises from geometric actions built via the method of co-adjoing orbits [24, 25, 26, 27, 28, 29, 30, 31, 32]. In [33], the method of co-adjoint orbits was used to produce the two-dimensional Wess-Zumino-Witten action [34, 35] and the Polyakov 2D quantum gravity action [36, 37]. More specifically, a vector field arises from the Kac-Moody algebra and is interpreted as a Yang-Mills field $A_{a}$, and a rank-two field $D_{ab}$ arises from the Virasoro algebra and is interpreted as the diffeomorphism field of TW gravity [33, 38, 39, 40, 41, 18, 19, 20, 42]. This interpretation becomes clear from the transformation properties of $A_{a}$ and $D_{ab}$ [33, 23, 42]: $A_{a}$ indeed transforms as a Yang-Mills field and $D_{ab}$ transforms as part of a projective connection of Thomas [1, 2].
In its original formulation [18, 19, 20] TW gravity was written in terms of the non-tensorial, but projectively invariant diffeomorphism field ${\mathcal{D}}_{ab}$, the non-affine but projectively invariant connection $\Pi^{c}{}_{ab}$, and the metric $g_{ab}$. The projective Schouten tensor ${\mathcal{P}}_{ab}$ was introduced in lieu of ${\mathcal{D}}_{ab}$, though this fails to be projectively invariant. Furthermore, the non-affine connection $\Pi^{c}{}_{ab}$ can be expanded in terms of the affine connection $\Gamma^{c}{}_{ab}$, though this connection is not by itself projectively invariant. Can we have the best of both worlds: a projectively invariant tensor field and a projectively invariant affine connection? The answer is yes, and the purpose of this paper is to demonstrate precisely how this is done. Expressing TW gravity in the resulting covariant and manifestly projectively invariant form is analogous to writing the Maxwell Lagrangian in a manifestly Lorentz invariant form using $F_{\mu\nu}$ rather than $\vec{E}$ and $\vec{B}$ fields. More to the point, a covariant and manifestly projective invariant formulation of TW gravity is analogous to the formulation of general relativity (GR) itself in terms of tensors and covariant derivatives rather than connections and parital derivatives.
As this new formulation of TW gravity is covariant, the usual tricks of GR become at our disposal: partial integration with respect to the covariant derivative is now available for instance. This was not available in the original formulation of TW gravity, and so deriving the field equations was quite cumbersome [18, 19, 20]. Now, with a covariant formulation, deriving the equations of motion becomes much more straightforward. Since the new formulation is manifestly projectively invariant, this simplifies how the action and field equations are expressed, much like, again, how GR is simpler to write in terms of the Ricci tensor and metric, rather than expanding fully in terms of derivatives of the metric.
We note a key distinction between the additional, projective direction of TW gravity and higher dimensions. In TW gravity, the projective direction can be thought of as a gauge dimension more so than an extra spacetime dimension. Upon integrating out this gauge dimension, TW gravity maintains its projective invariance on the ${\rm d}$-dimensional manifold $\mathcal{M}$, though it is no longer manifestly so when expressed in terms of the ${\rm d}$-dimensional projective Schouten tensor ${\mathcal{P}}_{ab}$, affine connection $\Gamma^{c}{}_{ab}$, and metric $g_{ab}$. The goal of this paper is to make projective invariance manifest on the ${\rm d}$-dimensional manifold $\mathcal{M}$ by introducing the projectively invariant Schouten tensor ${\bar{{\mathcal{P}}}}_{ab}$ and projectively invariant affine connection $\bar{\Gamma}^{c}{}_{ab}$ that is used to construct the projectively invariant covariant derivative ¯$\bar{\nabla}_{a}$. Thus TW gravity is effectively a ${\rm d}$-dimensional (set ${\rm d}=4$ to describe our Universe), projectively invariant model of gravity and not a higher dimensional gravitational model.
This paper is structured as follows. Section 2 is a review of TW gravity. Here the equations of motion are seen to be either manifestly projectively invariant or manifestly tensorial, but not both simultaneously. Section 3 contains the main results of the paper: the revelation of the covariant and manifestly projective invariant formulation of TW gravity. Here we present a covariant and manifestly projectively invariant derivation of the field equations, utilizing the familiar tensorial tricks of Einstein-Hilbert gravity such as integrating by parts with a covariant derivative. The independent fields we vary are the metric $g^{mn}$, the projectively invariant Schouten tensor ${\bar{{\mathcal{P}}}}_{ab}$, and the projectively invariant trace-removed Palatini tensor ${\tilde{C}}^{c}{}_{ab}$. The tensor ${\tilde{C}}^{c}{}_{ab}$ is the difference between the projectively invariant affine connection $\bar{\Gamma}^{c}{}_{ab}$ and the Levi-Civita connection $\hat{\Gamma}^{c}{}_{ab}$. We demonstrate the on-shell equivalence of the equations of motion across the different formulations in Section 4.
2 Review of TW Gravity
The foundations of TW gravity are with respect to a connection $\Gamma^{a}{}_{bc}$ that is not connected to a metric. In fact, to start there is no metric at all as TW gravity is based on the notion of equivalence of paths, regardless of any metric, as first investigated by Thomas [1, 2]. Transforming from one path to another path is known as a projective transformation in this context. This notion of projections is a generalization of the more familiar notion of projections in Euclidean geometry. In TW gravity, a projection indeed identifies points on one path with points on another path, but these points generally are not connected by straight lines as in a Euclidean projection. We define a projective transformation of a connection $\Gamma^{a}{}_{bc}$ as
$\Gamma^{a}{}_{bc}\to\Gamma^{a}{}_{bc}+\delta^{a}{}_{(b}v_{c)}$ (2.1)
with $v_{b}$ an arbitrary one form.
2.1 TW Connections
As originally presented in [18, 20], the TW connection $\tilde{\Gamma}^{\alpha}_{\ \beta\gamma}$ on the ${\rm d}+1$-dimensional Thomas cone $\mathcal{N}$ is expressed in terms of the fields $\mathcal{D}_{ab}$ and $\Pi^{a}{}_{bc}$ on the ${\rm d}$-dimensional base manifold $\mathcal{M}$
[…]
where
${\Pi}^{a}_{\,\,\,\,bc}=$ ${\Gamma}^{a}_{\,\,\,\,bc}+\delta^{a}{}_{(c}~{}\alpha_{b)}~{}~{}~{},~{}~{}~{}\alpha_{b}=-\tfrac{1}{{\rm d}+1}\Gamma^{c}{}_{bc}$ (2.3)
$\Upsilon^{\alpha}=$ $(0,0,\dots,0,\lambda)~{}~{}~{},~{}~{}~{}\alpha_{\beta}=\left(\alpha_{c},\lambda^{-1}\right)~{}~{}~{}.$ (2.4)
Latin indices $a,b,c,\dots$ are over the manifold $\mathcal{M}$ and Greek indices $\alpha,\beta,\gamma,\dots$ excluding $\lambda$ are over the manifold $\mathcal{N}$. The coordinates $x^{a}$ over $\mathcal{M}$ are extended to the coordinates $x^{\alpha}$ over $\mathcal{N}$ by appending the volume coordinate $\lambda$ as
$x^{\alpha}=$ $(x^{a},\lambda)~{}~{}~{}.$ (2.5)
The connection $\Gamma^{a}{}_{mn}$ transforms as an affine connection under general coordinate transformations on $\mathcal{M}$
${\Gamma}^{\prime a}{}_{mn}=$ $\frac{\partial x^{\prime a}}{\partial x^{b}}\frac{\partial x^{p}}{\partial x^{\prime m}}\frac{\partial x^{q}}{\partial x^{\prime n}}{\Gamma}^{b}{}_{pq}+\frac{\partial^{2}x^{b}}{\partial x^{\prime m}\partial^{\prime n}}\frac{\partial x^{\prime a}}{\partial x^{b}}~{}~{}~{}.$ (2.6)
Its normalized trace $\alpha_{b}$ therefore transforms on the manifold $\mathcal{M}$ as
$\alpha^{\prime}_{b}=$ $\frac{\partial x^{m}}{\partial x^{\prime b}}\alpha_{m}-\frac{1}{{\rm d}+1}\frac{\partial^{2}x^{m}}{\partial x^{\prime a}\partial x^{\prime b}}\frac{\partial x^{\prime a}}{\partial x^{m}}~{}~{}~{}.$ (2.7)
[…]
$x^{\prime\alpha}=$ $(x^{\prime 0}(x^{m}),x^{\prime 1}(x^{m}),\dots,x^{\prime{\rm d}-1}(x^{m}),\lambda^{\prime}=\lambda J^{1/({\rm d}+1)})~{}~{}~{},~{}~{}~{}J=|\partial x^{m}/\partial x^{\prime n}|$ (2.8)
the connection $\Gamma^{\alpha}{}_{\mu\nu}$ likewise transforms as an affine connection on $\mathcal{N}$
${\tilde{\Gamma}}^{\prime\alpha}{}_{\mu\nu}=$ $\frac{\partial x^{\prime\alpha}}{\partial x^{\rho}}\frac{\partial x^{\sigma}}{\partial x^{\prime\mu}}\frac{\partial x^{\beta}}{\partial x^{\prime\nu}}{\tilde{\Gamma}}^{\rho}{}_{\sigma\beta}+\frac{\partial^{2}x^{\beta}}{\partial x^{\prime\mu}\partial^{\prime\nu}}\frac{\partial x^{\prime\alpha}}{\partial x^{\beta}}~{}~{}~{}.$ (2.9)
[…]
$\alpha^{\prime}_{\beta}=$ $\frac{\partial x^{\mu}}{\partial x^{\prime\beta}}\alpha_{\mu}=-\frac{1}{{\rm d}+1}\delta^{b}{}_{\beta}\left(\frac{\partial x^{m}}{\partial x^{\prime b}}\Gamma^{n}{}_{mn}+\frac{\partial x^{\prime a}}{\partial x^{m}}\frac{\partial^{2}x^{m}}{\partial x^{\prime a}\partial x^{\prime b}}\right)+\delta^{\lambda}{}_{\beta}\lambda^{\prime-1}~{}~{}~{}.$ (2.10)
This is consistent with $\alpha_{b}$ transforming as the normalized trace of an affine connection on the manifold $\mathcal{M}$, Eq. (2.7), as demonstrated by the expansion following the second equality.
The fields $\mathcal{D}_{ab}$ and $\Pi^{a}{}_{bc}$ have the advantage of being projectively invariant under a projective transformation as defined in Eq. (2.1). The drawback of $\mathcal{D}_{ab}$ is that it is not a tensor and the drawback of $\Pi^{a}{}_{bc}$ is that it is not an affine connection. Instead, these transform under general coordinate transformations on $\mathcal{M}$ as
$\mathcal{D}^{\prime}_{ab}=$ $\frac{\partial x^{m}}{\partial x^{\prime a}}\frac{\partial x^{n}}{\partial x^{\prime b}}\left[\mathcal{D}_{mn}-\partial_{m}j_{n}-j_{m}j_{n}+j_{c}\Pi^{c}{}_{mn}\right]\;,$ (2.11)
${\Pi^{\prime}}^{a}_{\ bc}=$ $\frac{\partial x^{\prime a}}{\partial x^{p}}\frac{\partial x^{m}}{\partial x^{\prime b}}\frac{\partial x^{n}}{\partial x^{\prime c}}\Pi^{p}_{\ mn}+\frac{\partial x^{\prime a}}{\partial x^{m}}\frac{\partial^{2}x^{m}}{\partial x^{\prime b}\ \partial x^{\prime c}}-\delta^{a}_{\ (b}\frac{\partial x^{m}}{\partial x^{\prime c)}}j_{m}$ (2.12)
$j_{m}\equiv$ $\tfrac{1}{{\rm d}+1}\partial_{m}\log J=\tfrac{1}{{\rm d}+1}\frac{\partial x^{\prime q}}{\partial x^{p}}\frac{\partial^{2}x^{p}}{\partial x^{m}\partial x^{\prime q}}$ (2.13)
where we see the last three terms are the non-tensorial part of the $\mathcal{D}_{ab}$ transformation and the last term is the non-affine part of the $\Pi^{a}{}_{bc}$ transformation.
In particular, this means that the resulting gravitational model written in terms of $\mathcal{D}_{ab}$ and $\Pi^{a}{}_{bc}$ known as Thomas-Whitehead gravity [20] though manifestly projectively invariant does not have field equations that are manifestly tensorial, or even tensorial at all in the case of many of the field equations. The main goal of this paper is to express Thomas-Whitehead gravity in terms of a set of fields that are both manifestly tensorial and manifestly projectively invariant on the manifold $\mathcal{M}$. This will lead to field equations that are likewise manifestly tensorial and manifestly projectively invariant on the manifold $\mathcal{M}$, while at the same time being dynamically equivalent to the original set of field equations derived in [20] at least classically.
Derivative operators associated with each connection are defined as
~$\tilde{\nabla}_{\alpha}v_{\beta}=$ $\partial_{\alpha}v_{\beta}-\tilde{\Gamma}^{\mu}{}_{\alpha\beta}v_{\mu}~{}~{}~{},~{}~{}~{}\tilde{\nabla}_{\alpha}v^{\beta}=\partial_{\alpha}v_{\mu}+\tilde{\Gamma}^{\beta}{}_{\alpha\mu}v^{\mu}$ (2.14)
$\nabla_{a}v_{b}=$ $\partial_{a}v_{b}-\Gamma^{m}{}_{ab}v_{m}~{}~{}~{},~{}~{}~{}\nabla_{a}v^{b}=\partial_{\alpha}v_{\mu}+\Gamma^{b}{}_{am}v^{m}$ (2.15)
˘$\breve{\nabla}_{a}v_{b}=$ $\partial_{a}v_{b}-\Pi^{m}{}_{ab}v_{m}~{}~{}~{},~{}~{}~{}\breve{\nabla}_{a}v^{b}=\partial_{\alpha}v_{\mu}+\Pi^{b}{}_{am}v^{m}$ (2.16)
and similarly when acting on higher rank objects. The derivatives ~$\tilde{\nabla}_{\alpha}$ and $\nabla_{a}$ will be covariant derivatives when acting on tensors associated with their respective manifolds, as their associated connections are each affine on their respective manifolds. The derivative operator ˘$\breve{\nabla}_{a}$, however, will not generally be a covariant derivative regardless on what it acts as its associated connection $\Pi^{m}{}_{ab}$ is not affine.
[…]
$\mathcal{P}_{ab}$ $=\mathcal{D}_{ab}-\partial_{a}\alpha_{b}+\Gamma^{c}_{\ ab}\alpha_{c}+\alpha_{a}\alpha_{b}~{}~{}~{}.$ (2.17)
The tensor ${\mathcal{P}}_{ab}$, however, is not projectively invariant owing to its dependence on the connection $\Gamma^{c}{}_{ab}$.
In this paper, we will introduce the projectively invariant covariant derivative ¯$\bar{\nabla}_{a}$ (with connection $\bar{\Gamma}^{c}{}_{ab}$) and projectively invariant Schouten tensor ${\bar{{\mathcal{P}}}}_{ab}$, both being tensorial on $\mathcal{M}$ more specifically as well as projectively invariant. We then demonstrate that TW-gravity can be written in terms of these tensorial and projectively invariant ¯$\bar{\nabla}_{a}$ and ${\bar{{\mathcal{P}}}}_{ab}$. With this we will accomplish our goal of expressing TW-gravity in a covariant and manifestly projective invariant form in terms of ${\rm d}$-dimensional objects. At the same time, we will find the dynamics to be classically equivalent to the dynamics of TW-gravity written in terms of the original ˘$\breve{\nabla}_{a}$ and $\mathcal{D}_{ab}$.
2.2 TW Curvature Tensors
Curvature on the Thomas cone is defined in the usual way, through the commutator
[~$[\tilde{\nabla}_{\mu},\tilde{\nabla}_{\nu}]\mathcal{V}^{\alpha}\equiv\mathcal{K}^{\alpha}{}_{\beta\mu\nu}\mathcal{V}^{\beta}~{}~{}~{},~{}~{}~{}[\tilde{\nabla}_{\mu},\tilde{\nabla}_{\nu}]\mathcal{V}_{\beta}\equiv-\mathcal{K}^{\alpha}{}_{\beta\mu\nu}\mathcal{V}_{\alpha}$ (2.18)
with $\mathcal{V}^{\alpha}$ an arbitrary vector. This leads to the usual formula in terms of the connection
$\mathcal{K}^{\alpha}{}_{\beta\mu\nu}=$ $\partial_{\mu}\tilde{\Gamma}^{\alpha}{}_{\beta\nu}-\partial_{\nu}\tilde{\Gamma}^{\alpha}{}_{\beta\mu}+\tilde{\Gamma}^{\rho}_{\ \beta\nu}\tilde{\Gamma}^{\alpha}{}_{\mu\rho}-\tilde{\Gamma}^{\rho}{}_{\beta\mu}\tilde{\Gamma}^{\alpha}{}_{\nu\rho}~{}~{}~{}.$ (2.19)
We define an analogue of the Riemann tensor known as the equi-projective curvature tensor using $\Pi^{c}{}_{ab}$ as
$\mathcal{R}^{a}_{\ bcd}=\partial_{c}\Pi^{a}_{\ bd}-\partial_{d}\Pi^{a}_{\ bc}+\Pi^{e}_{\ bd}\Pi^{a}_{\ ce}-\Pi^{e}_{\ bc}\Pi^{a}_{\ de}$ (2.20)
which is projectively invariant by construction, but not tensorial. We can also define the Riemann curvature tensor, which will not generally be built from the Levi-Civita connection, and is given by
$R^{a}_{\ bcd}=\partial_{c}\Gamma^{a}_{\ bd}-\partial_{d}\Gamma^{a}_{\ bc}+\Gamma^{e}_{\ bd}\Gamma^{a}_{\ ce}-\Gamma^{e}_{\ bc}\Gamma^{a}_{\ de}~{}~{}~{}.$ (2.21)
Each of ${\mathcal{K}}^{a}{}_{bcd}$, $R^{a}{}_{bcd}$, and ${\mathcal{R}}^{a}{}_{bcd}$ are antisymmetric in their last two indices by construction
${\mathcal{K}}^{a}{}_{bcd}=-{\mathcal{K}}^{a}{}_{bdc}~{}~{}~{},~{}~{}~{}R^{a}{}_{bcd}=-R^{a}{}_{bdc}~{}~{}~{},~{}~{}~{}\mathcal{R}^{a}{}_{bcd}=-{\mathcal{R}}^{a}{}_{bdc}~{}~{}~{}.$ (2.22)
We form the Ricci tensor and its projectively invariant analogue on both manifold $\mathcal{M}$ and $\mathcal{N}$ as
$\mathcal{R}_{bd}\equiv$ $\mathcal{R}^{a}{}_{bad}$ (2.23)
$R_{bd}\equiv$ $R^{a}{}_{bad}$ (2.24)
$\mathcal{K}_{\beta\nu}=$ $\mathcal{K}^{\alpha}{}_{\beta\alpha\nu}~{}~{}~{}.$ (2.25)
It should be noted that both $\mathcal{R}_{bd}$ and $\mathcal{K}_{\beta\nu}$ are symmetric while $R_{bd}$ is not. The antisymmetric part of $R_{bd}$ is proportional to the curl of $\alpha_{b}$ as well as the antisymmetric part of $\mathcal{P}_{bd}$
$\mathcal{P}_{[bd]}=\frac{1}{d+1}R_{[bd]}=-\partial_{[b}\alpha_{d]}~{}~{}~{}.$ (2.26)
It is notable that the only non-vanishing components of $\mathcal{K}_{\beta\nu}$ are those on the manifold $\mathcal{M}$, namely $\mathcal{K}_{bn}$. To form scalars such as the Ricci tensor requires a metric, to which we turn to next.
2.3 The TW Action
Much of this section is a review of [20]. To construct an action from $\mathcal{K}^{\alpha}{}_{\beta\mu\nu}$, we need the metric on the Thomas cone
$G_{\alpha\beta}$ $=\delta^{a}_{\,\,\alpha}\delta^{b}_{\,\,\beta}\,g_{ab}-\lambda_{0}^{2}g_{\alpha}g_{\beta}$ (2.27)
$G^{\alpha\beta}$ $=g^{ab}(\delta^{\alpha}_{\ a}-g_{a}\Upsilon^{\alpha})(\delta^{\beta}_{\ b}-g_{b}\Upsilon^{\alpha})-\lambda_{0}^{-2}\Upsilon^{\alpha}\Upsilon^{\beta}~{}~{}~{}.$ (2.28)
The metric $g_{ab}$ is the metric on the ${\rm d}$-dimensional manifold $M$. The object $g_{\alpha}$ is defined as
$g_{\alpha}=(g_{a},\lambda^{-1})$ (2.29)
with $g_{a}$ defined as
$g_{a}=$ $-\frac{1}{d+1}\partial_{a}\log\sqrt{|g|}=-\tfrac{1}{{\rm d}+1}\hat{\Gamma}^{c}{}_{ac}$ (2.30)
$\hat{\Gamma}^{c}{}_{ab}=$ $\frac{1}{2}g^{cm}\left(\partial_{(a}g_{b)m}-\partial_{m}g_{ab}\right)~{}~{}~{}.$ (2.31)
[…]
$\Gamma^{c}{}_{ab}=$ $\hat{\Gamma}^{c}{}_{ab}+C^{c}{}_{ab}~{}~{}~{}.$ (2.32)
These relations to the Levi Civita connection will facilitate integration by parts when deriving the equations of motion.
Using Eq. (2.31) we can expand the Riemann curvature tensor as
$R^{a}_{\ bcd}=$ $\hat{R}^{a}_{\ bcd}+Q^{a}_{\ bcd}$ (2.33)
$Q^{a}_{\ bcd}=$$\hat{\nabla}_{c}C^{a}_{\ bd}-\hat{\nabla}_{d}C^{a}_{\ bc}+C^{e}_{\ bd}C^{a}_{\ ce}-C^{e}_{\ bc}C^{a}_{\ de}$ (2.34)
where $\hat{R}^{a}_{\ bcd}$ is the usual Levi-Civita Riemann tensor, built from $\hat{\Gamma}^{a}{}_{bc}$ rather than $\Gamma^{a}{}_{bc}$ as in Eq. (2.21), and ^$\hat{\nabla}_{a}$ the Levi-Cita covariant derivative for the connection $\hat{\Gamma}^{c}{}_{ab}$. The Levi-Civita Ricci curvature and scalar along with the analogous versions involving $Q^{a}{}_{bcd}$ are defined in the usual way
$\hat{R}_{bd}=$ $\hat{R}^{a}{}_{bad}~{}~{}~{},~{}~{}~{}\hat{R}=g^{bd}\hat{R}_{bd}~{}~{}~{}$ (2.35)
$Q_{bd}=$ $Q^{a}{}_{bad}~{}~{}~{},~{}~{}~{}Q=g^{bd}Q_{bd}~{}~{}~{}.$ (2.36)
With the metric $G_{\alpha\beta}$ now in hand from Eq. (2.27), we can construct invariants on the Thomas cone such as
$\mathcal{K}^{\alpha}{}_{\beta\mu\nu}\mathcal{K}_{\alpha}{}^{\beta\mu\nu}\equiv$ $\mathcal{K}^{\alpha}{}_{\beta\mu\nu}\mathcal{K}^{\rho}{}_{\sigma\kappa\tau}G_{\alpha\rho}G^{\beta\sigma}G^{\mu\kappa}G^{\nu\tau}$ (2.37)
$\mathcal{K}_{\alpha\beta}\mathcal{K}^{\alpha\beta}\equiv$ $\mathcal{K}^{\rho}{}_{\alpha\rho\beta}\mathcal{K}^{\sigma}{}_{\mu\sigma\nu}G^{\alpha\mu}G^{\beta\nu}$ (2.38)
$\mathcal{K}\equiv$ $\mathcal{K}_{\alpha\beta}G^{\alpha\beta}~{}~{}~{}.$ (2.39)
[…]
$K_{bcd}\equiv$ $g_{\alpha}\mathcal{K}^{\alpha}{}_{bcd}$ (2.40)
$=$ ˘$\breve{\nabla}_{[c}\mathcal{D}_{d]b}+g_{a}\mathcal{K}^{a}{}_{bcd}$ (2.41)
$=$ $\nabla_{[c}\mathcal{P}_{d]b}+\text{\textcent}_{a}\mathcal{K}^{a}{}_{bcd}~{}~{}~{}$ (2.42)
where ˘$\breve{\nabla}_{c}{\mathcal{D}}_{db}$ is defined as in Eq. (2.16) and ¢$\text{\textcent}_{a}$ is defined as
¢$\text{\textcent}_{a}\equiv$ $g_{a}-\alpha_{a}=\tfrac{1}{{\rm d}+1}C^{c}{}_{ac}~{}~{}~{}.$ (2.43)
The object ¢$\text{\textcent}_{a}$ is in fact a tensor on $\mathcal{M}$, as the non-tensorial parts of the transformations of $g_{a}$ and $\alpha_{a}$ cancel in the net transformation law for ¢$\text{\textcent}_{a}$.
The projective Cotton-York tensor is antisymmetric in its last two indices by construction and inheritance from the projective Riemann curvature tensor
$K_{bcd}=-K_{bdc}~{}~{}~{}.$ (2.44)
The projective Cotton-York tensor $K_{bcd}$ is clearly a tensor on both $\mathcal{M}$ and $\mathcal{N}$ as well as projectively invariant. In Eq. (2.40), the first line demonstrates the tensor character on $\mathcal{N}$ as recall both $g_{\alpha}$ and $\mathcal{K}^{\alpha}{}_{\beta\mu\nu}$ are tensors on $\mathcal{N}$. The second line demonstrates projective invariance even when restricted to the manifold $\mathcal{M}$ as each of $\mathcal{D}_{ab}$, $g_{a}$, and $\mathcal{K}^{a}{}_{bcd}$ are projectively invariant. The third line demonstrates the tensor character on the manifold $\mathcal{M}$ as each of $\mathcal{P}_{ab}$, ¢$\text{\textcent}_{a}$, and $\mathcal{K}^{a}{}_{bcd}$ are tensors on $\mathcal{M}$. None of the lines alone demonstrate both tensorality and projective invariance on the manifold $\mathcal{M}$, once again showing the importance of defining the tensorial and projectively invariant ${\bar{{\mathcal{P}}}}_{ab}$ and ¯$\bar{\nabla}_{a}$ that we will introduce later.
The TW action consists of two components, the projective Einstein-Hilbert action $S_{\text{PEH}}$ and the projective Gauss-Bonnet action $S_{\text{PGB}}$, which are given by
[…]
and
$S_{\text{PGB}}=-\frac{\tilde{J}_{0}c}{\lambda_{0}}\int d^{{\rm d}}xd\lambda\sqrt{|G|}\ \left(\mathcal{K}^{\alpha}{}_{\beta\mu\nu}\mathcal{K}_{\alpha}{}^{\beta\mu\nu}-4\mathcal{K}_{\alpha\beta}\mathcal{K}^{\alpha\beta}+\mathcal{K}^{2}\right)$ (2.46)
[…]
Our goal in this paper is to maintain this manifest covariance and projective invariance after reducing to the ${\rm d}$-dimensional manifold $\mathcal{M}$. To illuminate the behavior on the spacetime manifold $\mathcal{M}$ we can expand $S$ and integrate out $\lambda$
$S=$ $-\frac{1}{2\kappa_{0}}\int d^{\rm d}x\sqrt{|g|}\mathcal{K}+J_{0}c\lambda_{0}^{2}\int d^{\rm d}x\sqrt{|g|}K_{bcd}K^{bcd}$ (2.47)
$-J_{0}c\int d^{\rm d}x\sqrt{|g|}\left(\mathcal{K}_{abcd}\mathcal{K}^{abcd}-4\mathcal{K}_{bd}\mathcal{K}^{bd}+\mathcal{K}^{2}\right)~{}~{}~{}.$ (2.48)
The rescaled coupling constants are
[…]
A possible relationship between these scalings and renormalization in a Quantum Field Theory is under investigation.
On the manifold $\mathcal{M}$ we use the notation that indices are raised and lowered with $g^{ab}$ and $g_{ab}$, respectively
${\mathcal{K}}_{abcd}\equiv$ $g_{ma}{\mathcal{K}}^{m}{}_{bcd}$ (2.50)
${\mathcal{K}}^{abcd}=$ $g^{bp}g^{cm}g^{dn}{\mathcal{K}}^{a}{}_{pmn}$ (2.51)
$K^{bcd}\equiv$ $g^{bp}g^{cm}g^{dn}K_{pmn}~{}~{}~{}.$ (2.52)
On the manifold $\mathcal{N}$ whether indices are raised with $g^{ab}$ or $G^{\alpha\beta}$ is unambiguous for the projective Cotton-York tensor and the last three indices of the projective Riemann tensor $\mathcal{K}^{a}{}_{bcd}$. This is because $G^{ab}=g^{ab}$ and the lambda components of these tensor vanish aside from $\mathcal{K}^{\lambda}{}_{bcd}$. Care must be taken, however, to be clear what is meant by lowering the first index of $\mathcal{K}^{a}{}_{bcd}$. On the manifold $\mathcal{M}$, we no longer have the full $G_{\alpha\beta}$ as its $\lambda$ components and $\lambda$-dependence have been integrated out, and so raising and lowering is done via $g^{ab}$ and $g_{ab}$, respectively.
Quadratic dynamics for the fields can be seen in the projective Cotton-York tensor term when expanded using Eq. (2.40). The curvature tensors contain interactions as can be seen upon expanding in terms of the ${\rm d}$-dimensional $\mathcal{D}_{ab}$ and $\Pi^{c}{}_{ab}$ or ${\mathcal{P}}_{ab}$ and $\Gamma^{c}{}_{ab}$ as
$\mathcal{K}^{a}{}_{bcd}=$ $\mathcal{R}^{a}{}_{bcd}+\delta^{a}{}_{[c}\mathcal{D}_{d]b}=R^{a}{}_{bcd}+\delta^{a}{}_{[c}\mathcal{P}_{d]b}-\delta^{a}{}_{b}\mathcal{P}_{[cd]}$ (2.53)
$\mathcal{K}_{mn}=$ $\mathcal{R}_{mn}+(d-1)\mathcal{D}_{mn}=R_{mn}+(d-1)\mathcal{P}_{mn}-d\mathcal{P}_{[mn]}$ (2.54)
$\mathcal{K}=$ $\mathcal{R}+(d-1)\mathcal{D}=R+(d-1)\mathcal{P}~{}~{}~{}$ (2.55)
$\mathcal{R}\equiv$ $g^{ab}\mathcal{R}_{ab}~{}~{}~{},~{}~{}~{}\mathcal{D}\equiv g^{ab}\mathcal{D}_{ab}~{}~{}~{},~{}~{}~{}R\equiv g^{ab}R_{ab}~{}~{}~{},~{}~{}~{}{\mathcal{P}}\equiv g^{ab}{\mathcal{P}}_{ab}~{}~{}~{}.$ (2.56)
All other curvature components vanish aside from $\mathcal{K}^{\lambda}{}_{bcd}$ which can be solved for using the expressions of the projective Cotton-York tensor in Eq. (2.40). We see explicitly that TW gravity is not simultaneously manifestly both general coordinate invariant on $\mathcal{M}$ and projectively invariant when written in terms of these expansions. If the curvature tensors are expanded through the first equality in each of the above, they are manifestly projectively invariant, but not manifestly tensorial. If the second equality is used in each, they are no longer manifestly projectively invariant, but they are manifestly tensorial. The same is true for the projective Cotton-York tensor as was shown in Eq. (2.40). In the next section, we introduce the manifestly both tensorial and projectively invariant ${\bar{{\mathcal{P}}}}_{ab}$ and ¯$\bar{\nabla}_{a}$ and expand the above tensors in terms of these.
We conclude this section with a brief aside on the the values of the physical constants of TW gravity that are consistent with physical phenomena, specifically early Universe inflation. For ${\rm d}=4$ the units are
$\begin{split}[J_{0}]=\frac{ML^{2}}{T}\;,\quad[\mathcal{D}_{ab}]&=[R_{ab}]=L^{-2}\;,\quad\left[\kappa_{0}\right]=\frac{T^{2}}{ML}~{}~{}~{}.\cr\end{split}$ (2.57)
In this manner we see that $J_{0}$ has units of angular momentum, while $[\kappa_{0}]\propto\left[\frac{8\pi G}{c^{4}}\right]$ in accordance with General Relativity. Also, $\lambda$ and $\lambda_{0}$ both have units of length so their quotient is dimensionless. The parameter $\lambda_{0}$ sets the characteristic projective length scale appearing in the model. Using natural units $\hbar=c=1$ and defining the reduced Planck mass $M_{p}$ as
$M_{p}=\sqrt{\frac{\hbar c}{8\pi G}}$ (2.58)
a recent analysis in [23] suggests that if $\lambda_{0}\lesssim 6\cdot 10^{6}M_{p}^{-1}$ the inflaton $\phi\propto{\mathcal{P}}$ appearing naturally in TW gravity would have provided early Universe inflation matching current inflationary cosmology data within the $95\%$ confidence level. Specifically, it was found that to produce $N=60$ efolds, the parameters would have to satisfy the constraints $\lambda_{0}^{2}\lesssim 10^{3}J_{0}\kappa_{0}$ and $10^{8}\lesssim J_{0}\lesssim 4\cdot 10^{10}$ all within the $95\%$ confidence level. Furthermore, a value of $\kappa_{0}\sim M_{p}^{-2}$ would result in Einstein-Hilbert gravity + cosmological constant at the end of inflation within the $95\%$ confidence level  [^4] The value of $\kappa_{0}$ was not discussed in this way in [23]. However, it is easy to see that the TW action reduces to $(-2\kappa_{0})^{-1}(1+8\kappa_{0}J_{0}\lambda_{0}^{-1}\phi)R$ plus cosmological constant at the end of inflation if $\phi\propto{\mathcal{P}}$ settles to a constant value, with the traceless parts of $\mathcal{P}_{ab}$ equal to zero and assuming a Levi-Civita connection. The analysis in [23] leads to $18\kappa_{0}J_{0}\lambda_{0}^{-1}\phi<<$ 1 at the end of $N=60$ efolds of inflation within the $95\%$ confidence level. Thus the TW action reduces to approximately $(-2\kappa_{0})^{-1}R$ plus cosmological constant at the end of inflation, where we identify $\kappa_{0}\sim M_{p}^{-2}$ to match general relativity.. Putting this all together results in the above limit on $\lambda_{0}$. This suggest that the characteristic projective scale defined by $\lambda_{0}$ is less than a few million Planck lengths if it is TW gravity itself that is responsible for inflation. The importance of these results is that it appears inflation is automatically is built into the geometry of TW gravity with no additional field needed. The above analysis for the limit on $\lambda_{0}$ relies on the assumption that TW gravity must precisely match Einstein-Hilbert gravity + cosmological constant at the end of inflation to then transition to radiation dominance. Whether or not that is true depends on the specific interactions of TW gravity with matter because matter acts like radiation in the early Universe. These interactions are an ongoing avenue of research [18, 19, 45, 20].
2.4 Manifestly Projective Invariant form of the TW Field Equations
The equations of motion with respect to the independent fields ${\mathcal{D}}_{ab}$, $\Pi^{c}{}_{ab}$, and $g^{mn}$ were originally derived in [20]. The field equations for $\mathcal{D}_{ab}$ are  [^5] We have divided by a factor of $2\sqrt{|g|}$ from the original version presented in [20]. This makes the equation tensorial, though it is not manifest here.
$\mathbb{D}^{ab}\equiv-\tfrac{({\rm d}-1)}{4\kappa_{0}J_{0}}g^{ab}+\tfrac{1}{2}\hat{\mathcal{K}}_{c}^{\ (ab)c}+\tfrac{\lambda_{0}^{2}}{\sqrt{|g|}}\breve{\nabla}_{c}(\sqrt{|g|}K^{(ab)c})-\lambda_{0}^{2}g_{c}\breve{\mathcal{K}}^{(ab)c}=0$ (2.59)
where the following were introduced for convenience
$\hat{\mathcal{K}}_{d}^{\ abc}=$ $\mathcal{K}^{\bar{d}}_{\ \bar{a}\bar{b}\bar{c}}\mathcal{G}_{d\bar{d}}^{\ \ \ a\bar{a}[b|\bar{b}|c]\bar{c}}$ (2.60)
$\mathcal{G}_{d\bar{d}}^{\ \ \ a\bar{a}b\bar{b}c\bar{c}}=$ $G_{d\bar{d}}g^{a\bar{a}}g^{b\bar{b}}g^{c\bar{c}}-4\delta^{b}_{\ d}\delta^{\bar{b}}_{\ \bar{d}}g^{a\bar{a}}g^{c\bar{c}}+\delta^{b}_{\ d}\delta^{\bar{b}}_{\ \bar{d}}g^{ac}g^{\bar{a}\bar{c}}$ (2.61)
$\breve{\mathcal{K}}^{abc}=$ $g^{ap}g^{bm}g^{cn}\breve{\nabla}_{[m}{\mathcal{D}}_{n]p}~{}~{}~{}.$ (2.62)
The field equations for $\Pi^{c}{}_{ab}$ are  [^6] We have divided by a factor of $2\sqrt{|g|}$ from the original version presented in [20].
~$\tilde{\mathcal{E}}_{c}{}^{ab}\equiv\mathcal{E}_{c}^{\ ab}-\tfrac{1}{d+1}\delta_{c}^{\ (a}\mathcal{E}_{d}^{\ b)d}=0$ (2.63)
$\mathcal{E}_{c}^{\ ab}=\mathcal{E}_{c}^{\ ba}\equiv\tfrac{1}{4\kappa_{0}J_{0}\sqrt{|g|}}\breve{\nabla}_{c}(\sqrt{|g|}g^{ab})-\tfrac{1}{2\sqrt{|g|}}\breve{\nabla}_{d}(\sqrt{|g|}\hat{\mathcal{K}}_{c}^{\ (ab)d})\quad$
$\qquad\qquad\qquad\qquad+\tfrac{\lambda_{0}^{2}}{\sqrt{|g|}}\breve{\nabla}_{d}(\sqrt{|g|}g_{c}\breve{\mathcal{K}}^{(ab)d})-\lambda_{0}^{2}K^{(ab)d}\mathcal{D}_{dc}~{}~{}~{}.$ (2.64)
The field equations for $g^{mn}$ without any external matter sources are  [^7] This equation is equivalent to that originally presented in [20], though written in a slightly different way such that each individual term is projectively invariant.
$\mathcal{H}_{mn}=$ $-\mathcal{K}_{mn}-\kappa_{0}g_{mn}\left[\tfrac{2J_{0}\lambda_{0}^{2}}{({\rm d}+1)\sqrt{|g|}}\breve{\nabla}_{a}(\sqrt{|g|}{\mathcal{K}}^{a}{}_{bcd}K^{bcd})+\mathcal{L}\right]$ (2.65)
$+2\kappa_{0}J_{0}\lambda_{0}^{2}(K_{mcd}K_{n}{}^{cd}+2K_{bcm}K^{bc}{}_{n})+2\kappa_{0}J_{0}(8{\mathcal{K}}_{mb}{\mathcal{K}}^{b}{}_{n}-2{\mathcal{K}}{\mathcal{K}}_{mn})$ (2.66)
$+2\kappa_{0}J_{0}({\mathcal{K}}_{mbcd}{\mathcal{K}}_{n}{}^{bcd}-{\mathcal{K}}^{a}{}_{mcd}{\mathcal{K}}_{an}{}^{cd}-2{\mathcal{K}}^{abc}{}_{m}{\mathcal{K}}_{abcn})$ (2.67)
[…]
$\mathcal{L}=$ $-\tfrac{1}{2\kappa_{0}}{\mathcal{K}}+J_{0}\lambda_{0}^{2}K_{bcd}K^{bcd}-J_{0}(\mathcal{K}_{abcd}\mathcal{K}^{abcd}-4\mathcal{K}_{bd}\mathcal{K}^{bd}+\mathcal{K}^{2})~{}~{}~{}.$ (2.69)
All equations of motion are manifestly projectively invariant as they are composed entirely of projectively invariant objects $\mathcal{K}^{a}{}_{bcd}$, ˘$\breve{\nabla}_{a}$, $g_{ab}$, $g_{a}$, and $\mathcal{D}_{ab}$. None are manifestly tensorial, and in fact $\mathcal{E}_{c}{}^{ab}$ is not tensorial, though $\mathbb{D}^{ab}$ and $\mathcal{H}_{mn}$ are tensorial.
2.5 Covariant Form of the TW Field Equations
Rewriting the equations in terms of $\mathcal{P}_{ab}$, $\nabla_{a}$, and ¢$\text{\textcent}_{a}$ makes manifest the tensorial nature of $\mathbb{D}^{ab}$ and $\mathcal{H}_{mn}$ and illuminates how $\mathcal{E}_{c}{}^{ab}$ fails to be tensorial. The details of these recastings are found in appendix B. For $\mathbb{D}^{ab}$ we have
$\mathbb{D}^{ab}=$ $-\tfrac{d-1}{4J_{0}\kappa_{0}}g^{ab}-(d-1)g^{ab}\mathcal{K}+2(2d-3)\mathcal{K}^{ab}+\lambda_{0}^{2}(\nabla_{c}-({\rm d}+2)\text{\textcent}_{c})K^{(ab)c}$ (2.70)
which is manifestly tensorial for the reasons described above. To illustrate the tensorial nature of $\mathbb{H}_{mn}$, we express it in the familiar form of the Einstein Equations
$\hat{R}_{mn}-\tfrac{1}{2}g_{mn}\hat{R}=\kappa_{0}\hat{T}_{mn}$ (2.71)
$\hat{T}_{mn}=-\frac{1}{\kappa_{0}}\Big{(}Q_{mn}-\tfrac{1}{2}g_{mn}Q+({\rm d}-1)\mathcal{P}_{mn}-{\rm d}{\mathcal{P}}_{[mn]}-\tfrac{{\rm d}-1}{2}g_{mn}\mathcal{P}\Big{)}$
$\qquad+J_{0}g_{mn}\left[2\lambda_{0}^{2}(-\tfrac{1}{{\rm d}+1}\nabla_{a}+\text{\textcent}_{a})K_{bcd}\mathcal{K}^{abcd}-\lambda_{0}^{2}K^{bcd}K_{bcd}\right.$ (2.72)
$\left.\qquad\qquad\qquad+\mathcal{K}_{abcd}\mathcal{K}^{abcd}-4\mathcal{K}^{ab}\mathcal{K}_{ab}+\mathcal{K}^{2}\right]+2J_{0}\lambda_{0}^{2}(K_{mcd}K_{n}{}^{cd}+2K_{bcm}K^{bc}{}_{n})$
$\qquad+2J_{0}(8{\mathcal{K}}_{mb}{\mathcal{K}}^{b}{}_{n}-2{\mathcal{K}}{\mathcal{K}}_{mn})+2J_{0}({\mathcal{K}}_{mbcd}{\mathcal{K}}_{n}{}^{bcd}-{\mathcal{K}}^{a}{}_{mcd}{\mathcal{K}}_{an}{}^{cd}-2{\mathcal{K}}^{abc}{}_{m}{\mathcal{K}}_{abcn})$
where recall ${\mathcal{P}}_{mn}$, ${\mathcal{P}}$, $Q_{mn}$, and $Q$ were defined in Eqs. (2.17), (2.34), (2.36), and (2.56).
That $\mathcal{E}_{c}{}^{ab}$ is non-tensorial is easily seen when cast into the following form
$\mathcal{E}_{c}{}^{ab}=$ $\mathbb{E}_{c}{}^{ab}+g_{c}\mathbb{D}^{ab}+\delta^{(a}{}_{c}\mathfrak{g}^{b)}$ (2.73)
where for brevity we have defined
$\mathbb{E}_{c}{}^{ab}\equiv$ $(\nabla_{c}-2\text{\textcent}_{c})\mathscr{K}^{ab}-\delta^{(a}{}_{c}\text{\textcent}_{d}\mathscr{K}^{b)d}-[(\nabla_{d}-({\rm d}+1)\text{\textcent}_{d})\mathcal{K}_{c}{}^{(ab)d}-2\text{\textcent}_{c}\mathcal{K}^{ab}]$ (2.74)
$-\lambda_{0}^{2}(\mathcal{P}_{cd}-\nabla_{c}\text{\textcent}_{d}-\text{\textcent}_{c}\text{\textcent}_{d})K^{(ab)d}$ (2.75)
$\mathscr{K}^{ab}\equiv$ $(\tfrac{1}{4J_{0}\kappa_{0}}+\mathcal{K})g^{ab}-4\mathcal{K}^{ab}$ (2.76)
$\mathfrak{g}^{b}\equiv$ $\tfrac{1}{4J_{0}\kappa_{0}}g_{d}g^{bd}-\tfrac{1}{2}(\nabla_{d}-({\rm d}+3)\text{\textcent}_{d})(g^{bd}\mathcal{K}-4\mathcal{K}^{bd})~{}~{}~{}.$ (2.77)
The non-tensorality of $\mathcal{E}_{c}{}^{ab}$ is explicitly seen in the second and third terms of Eq. (2.73) along with the first term of Eq. (2.77): these take the form of the non-tensorial $g_{c}$ multiplying a tensor, the product of which is necessarily non-tensorial. An obvious question is: does the actual equation of motion $\tilde{E}_{c}{}^{ab}$, formed by subtracting from $\mathcal{E}_{c}{}^{ab}$ its trace, turn out to be tensorial? The answer is no as seen below
~$\tilde{\mathcal{E}}_{c}{}^{ab}=$ $\mathbb{E}_{c}{}^{ab}-\tfrac{1}{{\rm d}+1}\delta^{(a}{}_{c}\mathbb{E}_{m}{}^{b)m}+g_{c}\mathbb{D}^{ab}-\tfrac{1}{{\rm d}+1}\delta^{(a}{}_{c}g_{m}\mathbb{D}^{b)m}$ (2.78)
where we still see the presence of the non-tensorial $g_{c}$ times the tensor $\mathbb{D}^{ab}$. Since $\mathbb{E}_{c}{}^{ab}$ is itself a tensor, we do however see that ~$\tilde{\mathcal{E}}_{c}{}^{ab}$ becomes tensorial upon enforcing the equation of motion $\mathbb{D}^{ab}=0$.
Regardless, making the tensorial or non-tensorial nature of the equations manifest has now completely hidden their projective invariance. Now written in terms of the manifestly tensorial ${\mathcal{P}}_{ab}$, $\nabla_{a}$, ¢$\text{\textcent}_{a}$, and $Q_{ab}$, we have to expose the projective transformations on each of these, and show that they cancel, to demonstrate that the equations are indeed still projectively invariant. Under a projective transformation, Eq. (2.1), these transform as
$\mathcal{P}_{ab}\to$ $\mathcal{P}_{ab}+\nabla_{a}v_{b}-v_{a}v_{b}$ (2.79)
¢$\text{\textcent}_{a}\to$$\text{\textcent}_{a}+v_{b}$ (2.80)
$Q_{mn}\to$ $Q_{mn}+\nabla_{m}v_{n}-{\rm d}\nabla_{n}v_{m}+({\rm d}-1)v_{m}v_{n}~{}~{}~{}.$ (2.81)
In this venture, some helpful findings are that the following terms are all projectively invariant, though clearly not manifestly so
$\mathcal{P}_{cd}-\nabla_{c}\text{\textcent}_{d}-\text{\textcent}_{c}\text{\textcent}_{d}~{}~{}~{},~{}~{}~{}Q_{mn}-\tfrac{1}{2}g_{mn}Q+({\rm d}-1)\mathcal{P}_{mn}-{\rm d}{\mathcal{P}}_{[mn]}-\tfrac{{\rm d}-1}{2}g_{mn}\mathcal{P}$ (2.82)
$(-\tfrac{1}{{\rm d}+1}\nabla_{a}+\text{\textcent}_{a})K_{bcd}\mathcal{K}^{abcd}~{}~{}~{},~{}~{}~{}(\nabla_{c}-2\text{\textcent}_{c})\mathscr{K}^{ab}-\delta^{(a}{}_{c}\text{\textcent}_{d}\mathscr{K}^{b)d}$ (2.83)
$[(\nabla_{d}-({\rm d}+1)\text{\textcent}_{d})\mathcal{K}_{c}{}^{(ab)d}-2\text{\textcent}_{c}\mathcal{K}^{ab}]~{}~{}~{},~{}~{}~{}(\nabla_{c}-({\rm d}+2)\text{\textcent}_{c})K^{(ab)c}$ (2.84)
$(\nabla_{d}-({\rm d}+3)\text{\textcent}_{d})(g^{bd}\mathcal{K}-4\mathcal{K}^{bd})~{}~{}~{}.$ (2.85)
This section and the previous section have served to illustrate the main driving force for the introduction of manifestly tensorial and projectively invariant fields to which we now turn.
3 Covariant and Manifestly Projective Invariant
Formulation of TW Gravity
3.1 Projectively Invariant Connection and Schouten Tensor
We introduce a set of fields, denoted with overbars, that are both tensorial and projectively invariant. We begin by introducing a modification of the connection on the manifold $\mathcal{M}$ and define
$\bar{\Gamma}^{a}{}_{bc}$ $=\Gamma^{a}{}_{bc}-\delta^{a}{}_{(b}\text{\textcent}_{c)}=\Pi^{a}_{\ bc}-\delta^{a}_{\ (b}g_{c)}$ (3.1)
where ¢$\text{\textcent}_{c}$ and $g_{c}$ were defined in Eqs. (2.43) and (2.30). Modifying $\Pi^{a}{}_{bc}$ by subtracting the $g_{c}$ term has promoted the connection to affine. At the same time, changing $\Gamma^{a}{}_{bc}$ by subtracting the tensor ¢$\text{\textcent}_{c}$ has not changed its affine character. So $\bar{\Gamma}^{a}{}_{bc}$ indeed transforms as an affine connection, as in Eq. (2.6).
From $\bar{\Gamma}^{a}{}_{bc}$ we can construct a covariant derivative operator ¯$\bar{\nabla}_{a}$, which is manifestly tensorial from the first equality in Eq (3.1) and manifestly projectively invariant from the second. The operation of ¯$\bar{\nabla}_{a}$ is the usual rule for covariant derivatives on a rank-m covariant, rank-n contravariant arbitrary tensor $T_{a_{1}\dots a_{{\rm m}}}{}^{b_{1}\dots b_{{\rm n}}}$
¯$\bar{\nabla}_{c}T_{a_{1}\dots a_{{\rm m}}}{}^{b_{1}\dots b_{{\rm n}}}=\partial_{c}T_{a_{1}\dots a_{{\rm m}}}{}^{b_{1}\dots b_{{\rm n}}}$ $-\sum_{{\rm p}=1}^{{\rm m}}\bar{\Gamma}^{d}{}_{ca_{{\rm p}}}T_{a_{1}\dots a_{{\rm p}-1}da_{{\rm p}+1}\dots a_{{\rm m}}}{}^{b_{1}\dots b_{{\rm n}}}$ (3.2)
$+\sum_{{\rm p}=1}^{{\rm n}}\bar{\Gamma}^{b_{{\rm p}}}{}_{cd}T_{a_{1}\dots a_{{\rm m}}}{}^{b_{1}\dots b_{{\rm p}-1}db_{{\rm p}+1}\dots b_{{\rm n}}}~{}~{}~{}.$ (3.3)
Thus, substituting the defining Eq. (3.1), we have the following relationship between the three derivatives
¯$\bar{\nabla}_{c}T_{a_{1}\dots a_{{\rm m}}}{}^{b_{1}\dots b_{{\rm n}}}=(\nabla_{c}+({\rm m}-{\rm n})\text{\textcent}_{c})T_{a_{1}\dots a_{{\rm m}}}{}^{b_{1}\dots b_{{\rm n}}}$ $+\sum_{{\rm p}=1}^{{\rm m}}\text{\textcent}_{a_{{\rm p}}}T_{a_{1}\dots a_{{\rm p}-1}ca_{{\rm p}+1}\dots a_{{\rm m}}}{}^{b_{1}\dots b_{{\rm n}}}$ (3.4)
$-\sum_{{\rm p}=1}^{{\rm n}}\delta^{b_{{\rm p}}}{}_{c}~{}\text{\textcent}_{d}T_{a_{1}\dots a_{{\rm m}}}{}^{b_{1}\dots b_{{\rm p}-1}db_{{\rm p}+1}\dots b_{{\rm n}}}$ (3.5)
=(˘$=(\breve{\nabla}_{c}+({\rm m}-{\rm n})g_{c})T_{a_{1}\dots a_{{\rm m}}}{}^{b_{1}\dots b_{{\rm n}}}$ $+\sum_{{\rm p}=1}^{{\rm m}}g_{a_{{\rm p}}}T_{a_{1}\dots a_{{\rm p}-1}ca_{{\rm p}+1}\dots a_{{\rm m}}}{}^{b_{1}\dots b_{{\rm n}}}$ (3.6)
$-\sum_{{\rm p}=1}^{{\rm n}}\delta^{b_{{\rm p}}}{}_{c}~{}g_{d}T_{a_{1}\dots a_{{\rm m}}}{}^{b_{1}\dots b_{{\rm p}-1}db_{{\rm p}+1}\dots b_{{\rm n}}}.$ (3.7)
With this, it is straightforward to show that the projectively invariant terms listed in (2.83), (2.84), and (2.85) can all be written manifestly so as total derivatives with ¯$\bar{\nabla}_{a}$.
In terms of the Levi-Civita connection $\hat{\Gamma}^{a}{}_{bc}$, $\bar{\Gamma}^{a}{}_{bc}$ can be expressed as
$\bar{\Gamma}^{a}{}_{bc}=$ $\hat{\Gamma}^{a}{}_{bc}+\tilde{C}^{a}{}_{bc}$ (3.8)
$\tilde{C}^{a}{}_{bc}\equiv$ $C^{a}{}_{bc}-\tfrac{1}{{\rm d}+1}\delta^{a}{}_{(b}C^{m}{}_{c)m}~{}~{}~{}.$ (3.9)
This will be an important identity in the coming sections when we derive the field equations as it allows us to use partial integration with ¯$\bar{\nabla}_{a}$ equivalently to ^$\hat{\nabla}_{a}$ their partial integrations differ by a factor of $\tilde{C}^{c}{}_{bc}=0$. This feature forces the traces of the Levi-Civita connection and the projectively invariant connection to be identical. Indeed, tracing Eq. (3.8) and substituting Eq. (2.30) we find that $g_{a}$ is proportional to both the trace of the Levi-Civita connection and the trace of the projectively invariant connection
$g_{a}$ $=-\tfrac{1}{{\rm d}+1}\hat{\Gamma}^{c}{}_{ac}=-\tfrac{1}{{\rm d}+1}\bar{\Gamma}^{c}{}_{ac}~{}~{}~{}.$ (3.10)
We define the tensorial and manifestly projectively invariant tensor $\bar{\mathcal{P}}_{mn}$ by
$\bar{\mathcal{P}}_{mn}=\mathcal{P}_{mn}-\nabla_{m}\text{\textcent}_{n}-\text{\textcent}_{m}\text{\textcent}_{n}=\mathcal{D}_{mn}-\breve{\nabla}_{m}g_{n}-g_{m}g_{n}$ (3.11)
which is tensorial from the first equality and projectively invariant from the second. Comparing this with the first of the terms listed in (2.82) we see it is indeed nothing more than ${\bar{{\mathcal{P}}}}_{cd}$. We construct the trace in the usual manner, namely
$\bar{\mathcal{P}}=$ $g^{mn}\bar{\mathcal{P}}_{mn}=\mathcal{P}-\nabla^{m}\text{\textcent}_{m}-\text{\textcent}_{m}\text{\textcent}^{m}~{}~{}~{}.$ (3.12)
The new field $\bar{\mathcal{P}}_{mn}=\bar{\mathcal{P}}_{nm}$ is symmetric by construction and contains the same degrees of freedom as $\mathcal{D}_{bc}$ and $\mathcal{P}_{bc}$.
3.2 Projectively Invariant Curvature Tensors
From the projectively invariant connection, we build a projectively invariant Riemann curvature tensor
$\bar{R}^{a}_{\ bcd}=$ $\partial_{c}\bar{\Gamma}^{a}{}_{bd}-\partial_{d}\bar{\Gamma}^{a}{}_{bc}+\bar{\Gamma}^{m}{}_{bd}\bar{\Gamma}^{a}{}_{cm}-\bar{\Gamma}^{m}{}_{bc}\bar{\Gamma}^{a}{}_{dm}$ (3.13)
$=$ $R^{a}{}_{bcd}+\delta^{a}{}_{[c}\nabla_{d]}\text{\textcent}_{b}+\delta^{a}{}_{[c}\text{\textcent}_{d]}\text{\textcent}_{b}-\delta^{a}{}_{b}\nabla_{[c}\text{\textcent}_{d]}$ (3.14)
$=$ $\mathcal{R}^{a}{}_{bcd}+\delta^{a}{}_{[c}\breve{\nabla}_{d]}g_{b}+\delta^{a}{}_{[c}g_{d]}g_{b}$ (3.15)
which is manifestly covariant and projectively invariant when expressed as in the first line. In contrast, it is only manifestly covariant when expressed as in the second line and only manifestly projectively invariant when expressed as in the third line. We will express the other curvature tensors in this section in the same way: manifestly covariant and projectively invariant in their first lines, manifestly only one or the other in each of their other lines. The derivative ˘$\breve{\nabla}_{d}g_{b}$ is defined as in Eq. (2.16). It is worth noting that all three of the curvature tensors are commutators of derivative operators, and so are antisymmetric in their last two indices. However, the lack of metric compatibility means they are not generally antisymmetric in the first two indices (i.e. $R_{abcd}\neq-R_{bacd}$). Proceeding in the usual manner, we also construct the associated Ricci tensors by contracting over the first and third indices
$\bar{R}_{bd}=\bar{R}_{db}=$ $\bar{R}^{a}{}_{bad}$
$\qquad$== $\tfrac{1}{2}R_{(bd)}+\tfrac{{\rm d}-1}{2}\nabla_{(b}\text{\textcent}_{d)}+({\rm d}-1)\text{\textcent}_{b}\text{\textcent}_{d}$
$=$ $\mathcal{R}_{bd}+({\rm d}-1)\breve{\nabla}_{b}g_{d}+({\rm d}-1)g_{b}g_{d}~{}~{}~{}.$ (3.16)
We then take the trace to construct the Ricci scalars
$\bar{R}=g^{bd}\bar{R}_{bd}=$ $R+({\rm d}-1)\nabla^{b}\text{\textcent}_{b}+({\rm d}-1)\text{\textcent}^{b}\text{\textcent}_{b}$
$=$ $\mathcal{R}+({\rm d}-1)\breve{\nabla}^{b}g_{b}+({\rm d}-1)g^{b}g_{b}$ (3.17)
where raising/lowering is defined as occurring outside of the derivative $\nabla^{b}\equiv g^{bd}\nabla_{d}$ and ˘$\breve{\nabla}^{b}\equiv g^{bd}\breve{\nabla}_{d}$ and $g^{b}\equiv g^{bd}g_{d}$.
With the above considerations in mind, we can express the projective Riemann tensor in terms of projectively invariant fields
$\mathcal{K}^{a}{}_{bcd}=$ $\bar{R}^{a}{}_{bcd}+\delta^{a}{}_{[c}\bar{\mathcal{P}}_{d]b}$
$=$ $R^{a}{}_{bcd}+\delta^{a}{}_{[c}\mathcal{P}_{d]b}-\delta^{a}{}_{b}\mathcal{P}_{[cd]}$
$=$ $\mathcal{R}^{a}{}_{bcd}+\delta^{a}{}_{[c}\mathcal{D}_{d]b}~{}~{}~{}.$ (3.18)
The projective Ricci tensor is built by contracting $\mathcal{K}^{a}_{\ bcd}$ over the first and third indices:
$\mathcal{K}_{mn}=$ $\bar{R}_{mn}+(d-1)\bar{\mathcal{P}}_{mn}$
$=$ $R_{mn}+(d-1)\mathcal{P}_{mn}-d\mathcal{P}_{[mn]}$
$=$ $\mathcal{R}_{mn}+(d-1)\mathcal{D}_{mn}$ (3.19)
which is necessarily symmetric. We also form the projective Ricci scalar by taking the trace as
$\mathcal{K}=g^{mn}\mathcal{K}_{mn}=$ $\bar{R}+(d-1)\bar{\mathcal{P}}$
$=$ $R+(d-1)\mathcal{P}$
$=$ $\mathcal{R}+(d-1)\mathcal{D}~{}~{}~{}.$ (3.20)
Finally, the projective Cotton-York tensor is now expressed simply as
$K_{bcd}=$$\bar{\nabla}_{[c}\bar{\mathcal{P}}_{d]b}$
$=$ $\nabla_{[c}\mathcal{P}_{d]b}+\text{\textcent}_{a}\mathcal{K}^{a}{}_{bcd}$
$=$ ˘$\breve{\nabla}_{[c}\mathcal{D}_{d]b}+g_{a}\mathcal{K}^{a}{}_{bcd}~{}~{}~{}.$ (3.21)
In summary, we see each of the above curvature tensors are manifestly covariant and tensor in their first lines, where they are each written entirely in terms of the barred objects.
3.3 Covariant and Manifestly Projective Invariant TW Field Equations
Here we derive the covariant and manifestly projective invariant field equations by varying the action with respect to $\bar{{\mathcal{P}}}_{ab}$, ${\tilde{C}}^{c}{}_{ab}$, and $g^{mn}$. First, we point out two helpful tricks. One is that the divergence of an arbitrary vector $A^{a}$ is the same whether it is taken with ¯${\bar{\nabla}}_{a}$ or ^${\hat{\nabla}}_{a}$. This owes to the vanishing of the trace of ${\tilde{C}}^{a}{}_{bc}$ as follows
¯${\bar{\nabla}}_{a}A^{a}=$ $\partial_{a}A^{a}+{\bar{\Gamma}}^{a}{}_{ba}A^{a}$ (3.22)
$=$ $\partial_{a}A^{a}+\hat{\Gamma}^{a}{}_{ba}A^{a}+{\tilde{C}}^{a}{}_{ba}A^{a}$ (3.23)
$=$ $\partial_{a}A^{a}+\hat{\Gamma}^{a}{}_{ba}A^{a}$ (3.24)
$=$ ^$\hat{\nabla}_{a}A^{a}~{}~{}~{}.$ (3.25)
Second, we have that a total derivative ¯$\bar{\nabla}_{a}$ of an arbitrary vector $A^{a}$ is a surface term
$\int d^{\rm d}x\sqrt{g}\bar{\nabla}_{a}A^{a}=$ $\int d^{\rm d}x\sqrt{g}\hat{\nabla}_{a}A^{a}$ (3.26)
[…]
$=$ $\int d^{\rm d}x\partial_{a}(\sqrt{g}A^{a})=\text{surface term}$ (3.29)
where we have used Eq. (2.30) in going from the second to third lines. This means we can use partial integration with ¯$\bar{\nabla}_{a}$ at the Lagrangian density level at no cost to the field equations. For instance given an arbitrary rank three tensor $T^{abc}$
$T^{abc}\bar{\nabla}_{a}\delta\bar{{\mathcal{P}}}_{bc}=$ (¯$-(\bar{\nabla}_{a}T^{abc})\delta\bar{{\mathcal{P}}}_{bc}$ (3.30)
up to a surface term. Thus, we can work entirely with the TW Lagrangian density
$\mathcal{L}=-\frac{1}{2\kappa_{0}}\mathcal{K}d^{\rm d}x+J_{0}c\lambda_{0}^{2}K_{bcd}K^{bcd}-J_{0}c\left(g_{am}\mathcal{K}^{a}_{\ bcd}\mathcal{K}^{mbcd}-4\mathcal{K}_{bd}\mathcal{K}^{bd}+\mathcal{K}^{2}\right)$ (3.31)
and need not write out the full action.
The connection ¯$\bar{\nabla}_{a}$ provides as much calculational functionality in TW gravity as the Levi-Civita connection ^$\hat{\nabla}_{a}$ does in conventional general relativity. This is a powerful new feature that the projectively invariant and tensorial ¯$\bar{\nabla}_{a}$ and ${\bar{{\mathcal{P}}}}_{ab}$ provide for TW gravity, moving the results of the previous work [20] closer to the language of more conventional language in general relativity. In the following, we will demonstrate the utility of this new calculation tool.
3.3.1 Field Equations for ${\bar{{\mathcal{P}}}}_{ab}$
Varying the Lagrangian with respect to ${\bar{{\mathcal{P}}}}_{ab}$ depends only on the following two variations
$\delta_{{\bar{{\mathcal{P}}}}}{\mathcal{K}}^{a}{}_{bcd}=$ $\delta^{a}{}_{[c}\delta{\bar{{\mathcal{P}}}}_{d]b}~{}~{}~{},~{}~{}~{}\delta_{{\bar{{\mathcal{P}}}}}K_{bcd}={\bar{\nabla}}_{[c}\delta{\bar{{\mathcal{P}}}}_{d]b}$ (3.32)
along with contractions. Considering only these variations, we have
$\delta_{{\bar{{\mathcal{P}}}}}\mathcal{L}=$ $-\tfrac{1}{2\kappa_{0}}g^{ab}\delta_{{\bar{{\mathcal{P}}}}}{\mathcal{K}}_{ab}+2J_{0}\lambda_{0}^{2}K^{bcd}\delta_{{\bar{{\mathcal{P}}}}}K_{bcd}$ (3.33)
$-2J_{0}\left(g_{am}{\mathcal{K}}^{mbcd}\delta_{{\bar{{\mathcal{P}}}}}{\mathcal{K}}^{a}{}_{bcd}+(-4{\mathcal{K}}^{ab}+{\mathcal{K}}g^{ab})\delta_{{\bar{{\mathcal{P}}}}}{\mathcal{K}}_{ab}\right)$ (3.34)
$=$ $-\tfrac{1}{2\kappa_{0}}g^{ab}({\rm d}-1)\delta{\bar{{\mathcal{P}}}}_{ab}+2J_{0}\lambda_{0}^{2}K^{bcd}{\bar{\nabla}}_{[c}\delta{\bar{{\mathcal{P}}}}_{d]b}$ (3.35)
$-2J_{0}(g_{am}{\mathcal{K}}^{mbcd}\delta^{a}{}_{[c}\delta{\bar{{\mathcal{P}}}}_{d]b}+(-4{\mathcal{K}}^{ab}+{\mathcal{K}}g^{ab})({\rm d}-1)\delta{\bar{{\mathcal{P}}}}_{ab})~{}~{}~{}.$ (3.36)
Integrating by parts in the second term, contracting indices in the third, and renaming and reordering indices yields
$\delta_{{\bar{{\mathcal{P}}}}}\mathcal{L}=$ $\big{(}-\tfrac{{\rm d}-1}{2\kappa_{0}}g^{ab}+2J_{0}\lambda_{0}^{2}{\bar{\nabla}}_{c}K^{(ab)c}$ (3.37)
$\hskip 20.0pt-2J_{0}({\rm d}-1){\mathcal{K}}g^{ab}+8J_{0}({\rm d}-1){\mathcal{K}}^{ab}-4J_{0}{\mathcal{K}}^{ab}\big{)}\delta{\bar{{\mathcal{P}}}}_{ab}$ (3.38)
where we have used the antisymmetry of the last two indices of $K^{abc}$ to reorder the indices. Simplifying we have
$\delta_{{\bar{{\mathcal{P}}}}}\mathcal{L}=$ $2J_{0}\mathbb{D}^{ab}\delta{\bar{{\mathcal{P}}}}_{ab}$ (3.39)
where the equation of motion for ${\bar{{\mathcal{P}}}}_{ab}$ is
$\mathbb{D}^{ab}=$ $-\frac{d-1}{4J_{0}c\kappa_{0}}g^{ab}-(d-1)g^{ab}\mathcal{K}+2(2d-3)\mathcal{K}^{ab}+\lambda_{0}^{2}\bar{\nabla}_{c}K^{(ab)c}=0~{}~{}~{}.$ (3.40)
Using Eq. (3.4) to expand the last term, it is straightforward to show that this is the exact same as Eq. (2.70) that was derived via variation with respect to ${\mathcal{D}}_{ab}$. In the form of Eq. (3.40), however, we see this equation of motion is manifestly covariant and projectively invariant as it is written entirely in terms of such objects.
3.3.2 Field Equations for ${\tilde{C}}^{c}{}_{ab}$
Similar to ${\bar{{\mathcal{P}}}}_{ab}$, the only variations that contribute to the field equations for ${\tilde{C}}^{c}{}_{ab}$ are
$\delta_{{\bar{\Gamma}}}{\mathcal{K}}^{a}{}_{bcd}=$${\bar{\nabla}}_{[c}\delta{\bar{\Gamma}}^{a}{}_{d]b}~{}~{}~{},~{}~{}~{}\delta_{{\bar{\Gamma}}}K_{bcd}=-\delta{\bar{\Gamma}}^{m}{}_{b[c}{\bar{{\mathcal{P}}}}_{d]m}$ (3.41)
we will work with ${\bar{\Gamma}}^{c}{}_{ab}$ for most of this calculation, expanding in terms of ${\tilde{C}}^{c}{}_{ab}$ at the very end, so that we can borrow this calculation when we derive the metric field equations. Our first step parallels that for ${\bar{{\mathcal{P}}}}_{ab}$
$\delta_{{\bar{\Gamma}}}\mathcal{L}=$ $-\tfrac{1}{2\kappa_{0}}g^{ab}\delta_{{\bar{\Gamma}}}{\mathcal{K}}_{ab}+2J_{0}\lambda_{0}^{2}K^{bcd}\delta_{{\bar{\Gamma}}}K_{bcd}$ (3.42)
$-2J_{0}\left(g_{am}{\mathcal{K}}^{mbcd}\delta_{{\bar{\Gamma}}}{\mathcal{K}}^{a}{}_{bcd}+(-4{\mathcal{K}}^{ab}+{\mathcal{K}}g^{ab})\delta_{{\bar{\Gamma}}}{\mathcal{K}}_{ab}\right)$ (3.43)
$=$ $-\tfrac{1}{2\kappa_{0}}g^{ab}{\bar{\nabla}}_{[c}\delta{\bar{\Gamma}}^{c}{}_{b]a}+2J_{0}\lambda_{0}^{2}K^{bcd}(-\delta{\bar{\Gamma}}^{a}{}_{b[c}{\bar{{\mathcal{P}}}}_{d]m})$ (3.44)
$-2J_{0}(g_{am}{\mathcal{K}}^{mbcd}{\bar{\nabla}}_{[c}\delta{\bar{\Gamma}}^{a}{}_{d]b}+(-4{\mathcal{K}}^{ab}+{\mathcal{K}}g^{ab}){\bar{\nabla}}_{[c}\delta{\bar{\Gamma}}^{c}{}_{b]a})~{}~{}~{}.$ (3.45)
Integrating by parts, relabeling some indices, and collecting like terms and simplifying results in
$\delta_{{\bar{\Gamma}}}\mathcal{L}=$ $\delta{\bar{\Gamma}}^{c}{}_{a[b}\left\{{\bar{\nabla}}_{c]}\left[\tfrac{1}{2\kappa_{0}}g^{ab}+2J_{0}\left(-4{\mathcal{K}}^{ab}+{\mathcal{K}}g^{ab}\right)\right]\right\}$ (3.46)
$+\tfrac{1}{2}\left[-4J_{0}\lambda_{0}^{2}K^{(ab)d}{\bar{{\mathcal{P}}}}_{dc}+4J_{0}{\bar{\nabla}}_{d}\left(g_{cm}{\mathcal{K}}^{m(b|d|a)}\right)\right]\delta{\bar{\Gamma}}^{c}{}_{ab}~{}~{}~{}.$ (3.47)
Introducing a Kronecker delta symbols in the first term to move the antisymmetrization of indices off of $\delta{\bar{\Gamma}}^{c}{}_{ab}$ and ¯${\bar{\nabla}}_{c}$, using the antisymmetry property ${\mathcal{K}}^{mbda}=-{\mathcal{K}}^{mbad}$ in the second term on the second line, and simplifying results in
$\delta_{{\bar{\Gamma}}}\mathcal{L}=$ $2J_{0}\bigg{[}{\bar{\nabla}}_{m}\left(\tfrac{1}{4J_{0}\kappa_{0}}g^{a[b}\delta_{c}{}^{m]}-4{\mathcal{K}}^{a[b}\delta_{c}{}^{m]}+{\mathcal{K}}g^{a[b}\delta_{c}{}^{m]}\right)$ (3.48)
$\hskip 70.0pt-\lambda_{0}^{2}{\bar{{\mathcal{P}}}}_{dc}K^{(ab)d}-{\bar{\nabla}}_{d}\left(g_{cm}{\mathcal{K}}^{m(ab)d}\right)\bigg{]}\delta{\bar{\Gamma}}^{c}{}_{ab}~{}~{}~{}.$ (3.49)
Defining
$\mathbb{E}_{c}{}^{ab}=$${\bar{\nabla}}_{c}\left[g^{ab}\left({\mathcal{K}}+\tfrac{1}{4J_{0}\kappa_{0}}\right)-4{\mathcal{K}}^{ab}\right]-{\bar{\nabla}}_{d}\left(g_{cm}{\mathcal{K}}^{m(ab)d}\right)-\lambda_{0}^{2}{\bar{{\mathcal{P}}}}_{dc}K^{(ab)d}$ (3.50)
the variation Eq. (3.48) can be cast in the following form
$\delta_{{\bar{\Gamma}}}\mathcal{L}=$ $2J_{0}\left[\mathbb{E}_{c}{}^{ab}-{\bar{\nabla}}_{m}\left(\tfrac{1}{4J_{0}\kappa_{0}}g^{am}-4{\mathcal{K}}^{am}+{\mathcal{K}}g^{am}\right)\delta_{c}{}^{b}\right]\delta{\bar{\Gamma}}^{c}{}_{ab}~{}~{}~{}.$ (3.51)
To facilitate the goal of simultaneously isolating the equations of motion for ${\tilde{C}}^{c}{}_{ab}$ whilst keeping track of the contributions to the equations of motion for the metric $g^{mn}$, we define the traceless combination
$\tilde{\mathbb{E}}_{c}{}^{ab}\equiv$ $\mathbb{E}_{c}{}^{ab}-\tfrac{1}{{\rm d}+1}\delta_{c}{}^{(a}\mathbb{E}_{m}{}^{b)m}$ (3.52)
to express Eq. (3.51) as
$\delta_{{\bar{\Gamma}}}\mathcal{L}=$ $2J_{0}\left\{\tilde{\mathbb{E}}_{c}{}^{ab}+\tfrac{1}{{\rm d}+1}\delta_{c}{}^{(a}\mathbb{E}_{m}{}^{b)m}k\right.$ (3.53)
$\qquad\left.-\tfrac{1}{2}{\bar{\nabla}}_{m}\left[\delta_{c}{}^{(a}g^{b)m}\left(\tfrac{1}{4J_{0}\kappa_{0}}+{\mathcal{K}}\right)-4\delta_{c}{}^{(a}{\mathcal{K}}^{b)m}\right]\right\}\delta{\bar{\Gamma}}^{c}{}_{ab}~{}~{}~{}.$ (3.54)
Using Eq. (3.8) we split the variation Eq. 3.53 into two parts
$\delta_{{\bar{\Gamma}}}\mathcal{L}=$ $\delta_{{\hat{\Gamma}}}\mathcal{L}+\delta_{{\tilde{C}}}\mathcal{L}$ (3.55)
$\delta_{{\hat{\Gamma}}}\mathcal{L}=$ $2J_{0}\left\{\tilde{\mathbb{E}}_{c}{}^{ab}+\tfrac{1}{{\rm d}+1}\delta_{c}{}^{(a}\mathbb{E}_{m}{}^{b)m}\right.$ (3.56)
$\qquad\left.-\tfrac{1}{2}{\bar{\nabla}}_{m}\left[\delta_{c}{}^{(a}g^{b)m}\left(\tfrac{1}{4J_{0}\kappa_{0}}+{\mathcal{K}}\right)-4\delta_{c}{}^{(a}{\mathcal{K}}^{b)m}\right]\right\}\delta{\hat{\Gamma}}^{c}{}_{ab}$ (3.57)
$\delta_{{\tilde{C}}}\mathcal{L}=$ $2J_{0}\tilde{\mathbb{E}}_{c}{}^{ab}\delta{\tilde{C}}^{c}{}_{ab}$ (3.58)
where we have exploited the tracelessness of ${\tilde{C}}^{c}{}_{ab}$ to simplify $\delta_{\tilde{C}}\mathcal{L}$ to a single term involving $\tilde{\mathbb{E}}_{c}{}^{ab}$. Thus Eq. (3.58) yields the equations of motion for ${\tilde{C}}^{c}{}_{ab}$ and Eq. (3.56) leads to a contribution to the metric $g^{mn}$ field equations. Summarizing, the equations of motion for ${\tilde{C}}^{c}{}_{ab}$ are
$\tilde{\mathbb{E}}_{c}{}^{ab}\equiv$ $\mathbb{E}_{c}{}^{ab}-\tfrac{1}{{\rm d}+1}\delta_{c}{}^{(a}\mathbb{E}_{m}{}^{b)m}=0$ (3.59)
$\mathbb{E}_{c}{}^{ab}=$${\bar{\nabla}}_{c}\left[g^{ab}\left({\mathcal{K}}+\tfrac{1}{4J_{0}\kappa_{0}}\right)-4{\mathcal{K}}^{ab}\right]-{\bar{\nabla}}_{d}\left(g_{cm}{\mathcal{K}}^{m(ab)d}\right)-\lambda_{0}^{2}{\bar{{\mathcal{P}}}}_{dc}K^{(ab)d}~{}~{}~{}.$ (3.60)
The equations of motion for ${\tilde{C}}^{c}{}_{ab}$ are manifestly covariant and projectively invariant as they are composed entirely of such fields. Where the equations of motion for ${\bar{{\mathcal{P}}}}_{ab}$ and ${\mathcal{D}}_{ab}$ are identical, Eq. (3.59) for ${\tilde{C}}^{c}{}_{ab}$ is not the same as the equations of motion for $\Pi^{c}{}_{ab}$, Eq. (2.63). Utilizing Eq. (3.4) to expand the ¯$\bar{\nabla}_{c}$ derivatives and the definition of ${\bar{{\mathcal{P}}}}_{dc}$, Eq. (3.11), it is straightforward to show that the form of $\mathbb{E}_{c}{}^{ab}$ in Eq. (3.60) is identical to the form in Eq. (2.74). Therefore, the equations of motion for ${\tilde{C}}^{c}{}_{ab}$ indeed are related to those for $\Pi^{c}{}_{ab}$ as in Eq. (2.78). We rewrite the relationship again here in slightly more compact form
$\tilde{\mathbb{E}}_{c}{}^{ab}=$$\tilde{\mathcal{E}}_{c}{}^{ab}-\left(g_{c}\mathbb{D}^{ab}-\tfrac{1}{{\rm d}+1}\delta_{c}{}^{(a}g_{m}\mathbb{D}^{b)m}\right)~{}~{}~{}.$ (3.61)
We see then that the equations of motion for ${\tilde{C}}^{c}{}_{ab}$ and $\Pi^{a}{}_{ab}$ are identical upon enforcing the diff field equation $\mathbb{D}^{ab}=0$.
3.3.3 Field Equations for $g^{mn}$
As as segue to the full equations of motion for $g^{mn}$, we continue with Eq. (3.56) to rewrite it in terms of a variation with respect to the metric $g^{mn}$. We start by contracting the Kronecker delta’s to write this as
$\delta_{{\hat{\Gamma}}}\mathcal{L}=$ $2J_{0}(\tilde{\mathbb{E}}_{c}{}^{ab}\delta{\hat{\Gamma}}^{c}{}_{ab}+V^{a}\delta{\hat{\Gamma}}^{m}{}_{am})$ (3.62)
$V^{a}\equiv$ $\tfrac{2}{{\rm d}+1}\mathbb{E}_{m}{}^{am}-{\bar{\nabla}}_{m}\left[g^{am}\left(\tfrac{1}{4J_{0}\kappa_{0}}+{\mathcal{K}}\right)-4{\mathcal{K}}^{am}\right]$ (3.63)
$=$ $\tfrac{1}{{\rm d}+1}{\bar{\nabla}}_{b}\mathbb{D}^{ab}-\tfrac{\lambda_{0}^{2}}{{\rm d}+1}\left({\bar{\nabla}}_{b}{\bar{\nabla}}_{c}K^{(ab)c}+2{\bar{{\mathcal{P}}}}_{db}K^{bad}\right)$ (3.64)
where the equality between the definition of $V^{a}$ and the expression its second line is proved in appendix C.1. As proved in appendix C.2, a variation with respect to the Levi-Civita connection can be expressed as a variation with respect to the metric as
$\sqrt{g}A_{c}{}^{ab}\delta\hat{\Gamma}^{c}{}_{ab}=$ $\sqrt{g}\tfrac{1}{2}\left[\hat{\nabla}_{a}(A_{(mn)}{}^{a}-A^{a}{}_{mn})\right]\delta g^{mn}$ (3.65)
$\sqrt{g}A^{a}\delta\hat{\Gamma}^{m}{}_{am}=$ $\sqrt{g}\tfrac{1}{2}g_{mn}(\bar{\nabla}_{a}A^{a})\delta g^{mn}$ (3.66)
up to total derivatives and where $A_{c}{}^{ab}=A_{c}{}^{ba}$ is an arbitrary tensor symmetric in its last two indices and $A^{a}$ is an arbitrary vector. Using this, Eq. (3.62) becomes
$\delta_{{\hat{\Gamma}}}\mathcal{L}=$ $J_{0}\left[g_{mn}{\hat{\nabla}}_{a}V^{a}+{\hat{\nabla}}_{a}(\tilde{\mathbb{E}}_{(mn)}{}^{a}-\tilde{\mathbb{E}}^{a}{}_{mn})\right]\delta g^{mn}~{}~{}~{}.$ (3.67)
Using Eqs. (3.2)  and (3.8) to switch to ¯${\bar{\nabla}}_{a}$ we have
$\delta_{{\hat{\Gamma}}}\mathcal{L}=$ $J_{0}\Big{[}g_{mn}{\bar{\nabla}}_{a}V^{a}+{\bar{\nabla}}_{a}(\tilde{\mathbb{E}}_{(mn)}{}^{a}-\tilde{\mathbb{E}}^{a}{}_{mn})$ (3.68)
$+{\tilde{C}}^{p}{}_{a(m}\tilde{\mathbb{E}}_{|p|n)}{}^{a}+{\tilde{C}}^{p}{}_{a(m}\tilde{\mathbb{E}}_{n)p}{}^{a}-{\tilde{C}}^{p}{}_{a(m}\tilde{\mathbb{E}}^{a}{}_{n)p}\Big{]}\delta g^{mn}~{}~{}~{}.$ (3.69)
Finally, using the identity below that is proved in appendix C.3
¯${\bar{\nabla}}_{a}V^{a}=\tfrac{1}{{\rm d}+1}{\bar{\nabla}}_{a}{\bar{\nabla}}_{b}\mathbb{D}^{ab}-\tfrac{\lambda_{0}^{2}}{{\rm d}+1}{\bar{\nabla}}_{a}\left({\mathcal{K}}^{a}{}_{bcd}K^{bcd}\right)$ (3.70)
we have
$\delta_{{\hat{\Gamma}}}\mathcal{L}=$ $J_{0}\Big{[}-\tfrac{\lambda_{0}^{2}}{{\rm d}+1}g_{mn}{\bar{\nabla}}_{a}\left({\mathcal{K}}^{a}{}_{bcd}K^{bcd}\right)+\tfrac{1}{{\rm d}+1}g_{mn}{\bar{\nabla}}_{a}{\bar{\nabla}}_{b}\mathbb{D}^{ab}+{\bar{\nabla}}_{a}(\tilde{\mathbb{E}}_{(mn)}{}^{a}-\tilde{\mathbb{E}}^{a}{}_{mn})$ (3.71)
$\hskip 50.0pt+{\tilde{C}}^{p}{}_{a(m}\tilde{\mathbb{E}}_{|p|n)}{}^{a}+{\tilde{C}}^{p}{}_{a(m}\tilde{\mathbb{E}}_{n)p}{}^{a}-{\tilde{C}}^{p}{}_{a(m}\tilde{\mathbb{E}}^{a}{}_{n)p}\Big{]}\delta g^{mn}~{}~{}~{}.$ (3.72)
The rest of the terms involving $\delta g^{mn}$ will be found by expanding the action, Eq. (3.31), to expose the factors of the inverse metric $g^{mn}$
$\mathcal{L}=$ $-\frac{1}{2\kappa_{0}}g^{mn}\mathcal{K}_{mn}+J_{0}\lambda_{0}^{2}g^{ab}g^{cm}g^{dn}K_{bcd}K_{amn}$ (3.73)
$-J_{0}\left(g_{am}g^{bn}g^{cp}g^{dq}\mathcal{K}^{a}_{\ bcd}\mathcal{K}^{m}{}_{npq}-4g^{am}g^{bn}\mathcal{K}_{ab}\mathcal{K}_{mn}+g^{ab}g^{mn}\mathcal{K}_{ab}{\mathcal{K}}_{mn}\right).$ (3.74)
Next, the variation begins by varying the product of the Lagrangian density with $\sqrt{g}$
$\delta_{g}(\sqrt{g}\mathcal{L})=$ $\tfrac{\sqrt{g}}{2}g^{mn}\delta g_{mn}\mathcal{L}-\tfrac{\sqrt{g}}{2\kappa_{0}}\delta g^{mn}{\mathcal{K}}_{mn}$ (3.75)
$+\sqrt{g}J_{0}\lambda_{0}^{2}(\delta g^{ab}K_{acd}K_{b}{}^{cd}+2\delta g^{dn}K_{bcd}K^{bc}{}_{n})$ (3.76)
$-\sqrt{g}J_{0}(\delta g_{am}{\mathcal{K}}^{a}{}_{bcd}{\mathcal{K}}^{mbcd}+\delta g^{bn}{\mathcal{K}}^{a}{}_{bcd}{\mathcal{K}}_{an}{}^{cd}+2\delta g^{dq}{\mathcal{K}}^{a}{}_{bcd}{\mathcal{K}}_{a}{}^{bc}{}_{q})$ (3.77)
$+\sqrt{g}J_{0}(8\delta g^{am}{\mathcal{K}}_{ab}{\mathcal{K}}_{m}{}^{b}+2\delta g^{mn}{\mathcal{K}}_{mn}{\mathcal{K}})+\sqrt{g}\delta_{\hat{\Gamma}}\mathcal{L}~{}~{}~{}.$ (3.78)
Recall here on the manifold $\mathcal{M}$, we raise and lower indices with $g^{mn}$ and $g_{mn}$ as for example ${\mathcal{K}}_{a}{}^{bc}{}_{q}=g_{am}g^{bn}g^{cp}{\mathcal{K}}^{m}{}_{npq}$. Using Eq. (3.71), collecting like terms, and relabeling indices we have
$\delta_{g}(\sqrt{g}\mathcal{L})=$ $\tfrac{\sqrt{g}}{2\kappa_{0}}\delta g^{mn}\bigg{\{}-{\mathcal{K}}_{mn}-\kappa_{0}g_{mn}\left[\tfrac{2J_{0}\lambda_{0}^{2}}{{\rm d}+1}{\bar{\nabla}}_{a}\left({\mathcal{K}}^{a}{}_{bcd}K^{bcd}\right)+\mathcal{L}\right]$ (3.79)
$+2\kappa_{0}J_{0}\lambda_{0}^{2}\left(K_{mcd}K_{n}{}^{cd}+2K_{bcm}K^{bc}{}_{n}\right)+2\kappa_{0}J_{0}\lambda_{0}^{2}\left(8{\mathcal{K}}_{mb}{\mathcal{K}}_{n}{}^{b}-2{\mathcal{K}}{\mathcal{K}}_{mn}\right)$ (3.80)
$+2\kappa_{0}J_{0}\left({\mathcal{K}}_{mbcd}{\mathcal{K}}_{n}{}^{bcd}-{\mathcal{K}}^{a}{}_{mcd}{\mathcal{K}}_{an}{}^{cd}-2{\mathcal{K}}^{abc}{}_{m}{\mathcal{K}}_{a}{}_{bcn}\right)+\tfrac{2\kappa_{0}J_{0}}{{\rm d}+1}g_{mn}{\bar{\nabla}}_{a}{\bar{\nabla}}_{b}\mathbb{D}^{ab}$ (3.81)
$+2\kappa_{0}J_{0}\left[{\bar{\nabla}}_{a}(\tilde{\mathbb{E}}_{(mn)}{}^{a}-\tilde{\mathbb{E}}^{a}{}_{mn})+{\tilde{C}}^{p}{}_{a(m}\tilde{\mathbb{E}}_{|p|n)}{}^{a}\right.$ (3.82)
$\hskip 100.0pt\left.+{\tilde{C}}^{p}{}_{a(m}\tilde{\mathbb{E}}_{n)p}{}^{a}-{\tilde{C}}^{p}{}_{a(m}\tilde{\mathbb{E}}^{a}{}_{n)p}\right]\bigg{\}}~{}~{}~{}.$ (3.83)
We collect the factor in the curly brackets in an equation of motion tensor $\mathbb{H}_{mn}$
$\delta_{g}(\sqrt{g}\mathcal{L})=$ $\tfrac{\sqrt{g}}{2\kappa_{0}}\delta g^{mn}\mathbb{H}_{mn}$ (3.84)
where the associated equation of motion is
$0=\mathbb{H}_{mn}=$ $\mathcal{H}_{mn}+\tfrac{2\kappa_{0}J_{0}}{{\rm d}+1}g_{mn}{\bar{\nabla}}_{a}{\bar{\nabla}}_{b}\mathbb{D}^{ab}+2\kappa_{0}J_{0}\left[{\bar{\nabla}}_{a}(\tilde{\mathbb{E}}_{(mn)}{}^{a}-\tilde{\mathbb{E}}^{a}{}_{mn})\right.$ (3.85)
$\left.\hskip 120.0pt+{\tilde{C}}^{p}{}_{a(m}\tilde{\mathbb{E}}_{|p|n)}{}^{a}+{\tilde{C}}^{p}{}_{a(m}\tilde{\mathbb{E}}_{n)p}{}^{a}-{\tilde{C}}^{p}{}_{a(m}\tilde{\mathbb{E}}^{a}{}_{n)p}\right]$ (3.86)
$\mathcal{H}_{mn}=$ $-\mathcal{K}_{mn}-\kappa_{0}g_{mn}\left[\tfrac{2J_{0}\lambda_{0}^{2}}{{\rm d}+1}{\bar{\nabla}}_{a}\left({\mathcal{K}}^{a}{}_{bcd}K^{bcd}\right)+\mathcal{L}\right]$ (3.87)
$+2\kappa_{0}J_{0}\lambda_{0}^{2}(K_{mcd}K_{n}{}^{cd}+2K_{bcm}K^{bc}{}_{n})+2\kappa_{0}J_{0}(8{\mathcal{K}}_{mb}{\mathcal{K}}^{b}{}_{n}-2{\mathcal{K}}{\mathcal{K}}_{mn})$ (3.88)
$+2\kappa_{0}J_{0}({\mathcal{K}}_{mbcd}{\mathcal{K}}_{n}{}^{bcd}-{\mathcal{K}}^{a}{}_{mcd}{\mathcal{K}}_{an}{}^{cd}-2{\mathcal{K}}^{abc}{}_{m}{\mathcal{K}}_{abcn})~{}~{}~{}.$ (3.89)
To show that $\mathcal{H}_{mn}$ is equivalent to the form in Eq. (2.65), Eq. (3.4) is used to relate the ¯${\bar{\nabla}}_{a}$ term to a ˘$\breve{\nabla}_{a}$ term as
¯${\bar{\nabla}}_{a}({\mathcal{K}}^{a}{}_{bcd}K^{bcd})=(\breve{\nabla}_{a}-({\rm d}+1)g_{a})({\mathcal{K}}^{a}{}_{bcd}K^{bcd})=(\nabla_{a}-({\rm d}+1)\text{\textcent}_{a})({\mathcal{K}}^{a}{}_{bcd}K^{bcd}).$ (3.90)
4 Classical Equivalence of Field Equations Across Formulations
We have seen that the equations of motion derived in terms of the old fields $\mathcal{D}_{ab}$, $\Pi^{a}{}_{bc}$, and $g^{mn}$ are equivalent to those derived in terms of the new fields ${\bar{{\mathcal{P}}}}_{ab}$, ${\tilde{C}}^{a}{}_{bc}$, and $g^{mn}$ upon enforcement of the equations of motion (on-shell). They describe the same classical system, regardless of field choice, as expected. Eq. (3.61) is a particularly interesting relationship between $\tilde{\mathbb{E}}_{c}{}^{ab}$ and ~$\tilde{\mathcal{E}}_{c}{}^{ab}$ and $g_{c}\mathbb{D}^{ab}$ as just the right combination to remove the non-tensorial part of ~$\tilde{\mathcal{E}}_{c}{}^{ab}$ to create the tensorial $\tilde{\mathbb{E}}_{c}{}^{ab}$. Furthermore, we see that the tensorial result $\tilde{\mathbb{E}}_{c}{}^{ab}$ is also manifestly projectively invariant as it is composed entirely of such objects ~$\tilde{\mathcal{E}}_{c}{}^{ab}$ and $g_{c}\mathbb{D}^{ab}$.
In the following, we investigate some nuances in the variation of the Lagrangian and summarize the field equations and their classical equivalence. We start by investigating the tensorial or non-tensorial character of the fields and their variations. We suspect that this will have interesting consequences at the quantum level in terms of Faddeev-Popov-like projective gauge fixing conditions. It is therefore important to work out the details at the classical level here first, leaving the quantum calculation for a future endeavor.
4.1 Equivalence of the Variation of the Action
Let us start by summarizing the variation in terms of the old fields
$\delta S=2J_{0}\int d^{\rm d}x\sqrt{g}\left(\mathbb{D}^{ab}\delta{\mathcal{D}}_{ab}+\tilde{\mathcal{E}}_{c}{}^{ab}\delta\Pi^{c}{}_{ab}+\tfrac{1}{4J_{0}\kappa_{0}}\mathcal{H}_{mn}\delta g^{mn}\right)~{}~{}~{}.$ (4.1)
Using Eqs. (3.11), (3.1), and (3.8) we can solve for the old field variations $\delta{\mathcal{D}}_{ab}$ and $\delta\Pi^{c}{}_{ab}$ in terms of the new field variations $\delta{\bar{{\mathcal{P}}}}_{ab}$ and $\delta{\tilde{C}}^{c}{}_{ab}$
$\delta\mathcal{D}_{ab}=$ $\delta\bar{\mathcal{P}}_{ab}-\tfrac{1}{{\rm d}+1}{\bar{\nabla}}_{a}\delta\hat{\Gamma}^{m}{}_{bm}-g_{c}(\delta\hat{\Gamma}^{c}{}_{ab}-\tfrac{1}{{\rm d}+1}\delta^{c}{}_{(a}\delta\hat{\Gamma}^{m}{}_{b)m})-g_{c}\delta{\tilde{C}}^{c}{}_{ab}$ (4.2)
$\delta\Pi^{c}{}_{ab}=$ $\delta\hat{\Gamma}^{c}{}_{ab}+\delta{\tilde{C}}^{c}{}_{ab}+\delta^{c}{}_{(a}\delta g_{b)}~{}~{}~{}.$ (4.3)
With this substitution, the variation becomes
$\delta S=2J_{0}\int d^{\rm d}x\sqrt{g}$ $\big{(}\mathbb{D}^{ab}\delta{\bar{{\mathcal{P}}}}_{ab}+\tilde{\mathcal{E}}_{c}{}^{ab}\delta{\tilde{C}}^{c}{}_{ab}-g_{c}\mathbb{D}^{ab}\delta{\tilde{C}}^{c}{}_{ab}+\tfrac{1}{4J_{0}\kappa_{0}}\mathcal{H}_{mn}\delta g^{mn}$ (4.4)
$-\tfrac{1}{{\rm d}+1}\mathbb{D}^{ab}{\bar{\nabla}}_{a}\delta\hat{\Gamma}^{m}{}_{bm}-g_{c}\mathbb{D}^{ab}(\delta\hat{\Gamma}^{c}{}_{ab}-\tfrac{1}{{\rm d}+1}\delta^{c}{}_{(a}\delta\hat{\Gamma}^{m}{}_{b)m})+\tilde{\mathcal{E}}_{c}{}^{ab}\delta\hat{\Gamma}^{c}{}_{ab}\big{)}.$ (4.5)
The third term on the first line and second term on the second line can be rewritten as follows
$g_{c}\mathbb{D}^{ab}\delta{\tilde{C}}^{c}{}_{ab}=$ $(g_{c}\mathbb{D}^{ab}-\tfrac{1}{{\rm d}+1}\delta^{(a}{}_{c}g_{m}\mathbb{D}^{b)m})\delta{\tilde{C}}^{c}{}_{ab}$ (4.6)
$g_{c}\mathbb{D}^{ab}(\delta\hat{\Gamma}^{c}{}_{ab}-\tfrac{1}{{\rm d}+1}\delta^{c}{}_{(a}\delta\hat{\Gamma}^{m}{}_{b)m})=$ $(g_{c}\mathbb{D}^{ab}-\tfrac{1}{{\rm d}+1}\delta^{(a}{}_{c}g_{m}\mathbb{D}^{b)m})\delta\hat{\Gamma}^{c}{}_{ab}~{}~{}~{}.$ (4.7)
Essentially, fully contracting a traceless object with another enforces tracelessness on both.
Substituting Eqs. (4.6) and (4.7) into Eq. (4.4), the second and third terms on the first line can be combined using Eq.  (3.61), as can the second and third terms on the second line. Furthermore, in Eq. (4.4) the first term on the second line can be integrated by parts. All this leaves us with
$\delta S=2J_{0}\int d^{\rm d}x\sqrt{g}$ $\big{(}\mathbb{D}^{ab}\delta{\bar{{\mathcal{P}}}}_{ab}+\tilde{\mathbb{E}}_{c}{}^{ab}\delta{\tilde{C}}^{c}{}_{ab}+\tfrac{1}{4J_{0}\kappa_{0}}\mathcal{H}_{mn}\delta g^{mn}$ (4.8)
$+\tfrac{1}{{\rm d}+1}({\bar{\nabla}}_{a}\mathbb{D}^{ab})\delta\hat{\Gamma}^{m}{}_{bm}+\tilde{\mathbb{E}}_{c}{}^{ab}\delta\hat{\Gamma}^{c}{}_{ab}\big{)}~{}~{}~{}.$ (4.9)
Using Eqs. (3.65) and (3.66) we can rewrite the second line in terms of the variation of $\delta g^{mn}$
$\delta S=2J_{0}\int d^{\rm d}x\sqrt{g}$ $\big{[}\mathbb{D}^{ab}\delta{\bar{{\mathcal{P}}}}_{ab}+\tilde{\mathbb{E}}_{c}{}^{ab}\delta{\tilde{C}}^{c}{}_{ab}+\tfrac{1}{4J_{0}\kappa_{0}}\mathcal{H}_{mn}\delta g^{mn}$ (4.10)
$+\left(\tfrac{1}{2({\rm d}+1)}g_{mn}{\bar{\nabla}}_{a}{\bar{\nabla}}_{b}\mathbb{D}^{ab}+\tfrac{1}{2}\hat{\nabla}_{a}(\tilde{\mathbb{E}}_{(mn)}{}^{a}-\tilde{\mathbb{E}}^{a}{}_{mn})\delta g^{mn}\right)\big{]}~{}.$ (4.11)
Switching from ^$\hat{\nabla}_{a}$ to ¯${\bar{\nabla}}_{a}$ via Eq. (3.8), we find that the resulting coefficients of all $\delta g^{mn}$ terms combine to form $\mathbb{H}_{mn}$ as defined in Eq. (3.85). This leaves us with our final result:
$\delta S=2J_{0}\int d^{\rm d}x\sqrt{g}\left(\mathbb{D}^{ab}\delta{\bar{{\mathcal{P}}}}_{ab}+\tilde{\mathbb{E}}_{c}{}^{ab}\delta{\tilde{C}}^{c}{}_{ab}+\tfrac{1}{4J_{0}\kappa_{0}}\mathbb{H}_{mn}\delta g^{mn}\right)~{}~{}~{}.$ (4.12)
All individual equations of motion and variations in Eq. (4.12) in terms of the new fields are tensorial. This is not true for Eq. (4.1) in terms of the old fields. It must be true, however, that the variation of the action is general coordinate invariant either way. What saves Eq. (4.1), maintaining its general coordinate invariance, is that the non-tensorial parts of $\delta{\mathcal{D}}_{ab}$ and ~$\tilde{\mathcal{E}}_{c}{}^{ab}$ conspire to cancel in Eq. (4.1) while all other fields and variations are tensorial. Specifically, performing a coordinate transformation on Eqs. (4.2) and (2.78) we find
$\delta{\mathcal{D}}^{\prime}_{ab}=$ $\frac{\partial x^{m}}{\partial x^{\prime a}}\frac{\partial x^{n}}{\partial x^{\prime b}}\left[\delta{\mathcal{D}}_{mn}+j_{c}\delta\Pi^{c}{}_{mn}\right]$ (4.13)
~$\tilde{\mathcal{E}}^{\prime}_{c}{}^{ab}=$ $\frac{\partial x^{p}}{\partial x^{\prime c}}\frac{\partial x^{\prime a}}{\partial x^{\prime m}}\frac{\partial x^{\prime b}}{\partial x^{n}}\left[\tilde{\mathcal{E}}_{p}{}^{mn}-j_{p}\mathbb{D}^{mn}+\tfrac{1}{{\rm d}+1}\delta_{p}{}^{(m}j_{d}\mathbb{D}^{n)d})\ \right]$ (4.14)
$j_{d}=$ $\tfrac{1}{{\rm d}+1}\partial_{m}\log J=\tfrac{1}{{\rm d}+1}\frac{\partial x^{\prime q}}{\partial x^{p}}\frac{\partial^{2}x^{p}}{\partial x^{m}\partial x^{\prime q}}~{}~{}~{},~{}~{}~{}J=|\partial x^{p}/\partial x^{\prime q}|$ (4.15)
where the non-tensorial parts of these transformations stem from their dependence on $g_{a}$ as all other parts are tensorial (though a connection itself is not tensorial, its variation is tensorial). Note that the above result for $\delta\mathcal{D}^{\prime}_{ab}$ is consistent with taking the functional variation of Eq. (2.11) as $\delta j_{m}=0$ necessarily.
With the above results for $\delta{\mathcal{D}}^{\prime}_{ab}$ and ~$\tilde{\mathcal{E}}^{\prime}_{c}{}^{ab}$, we find that the following combination found in Eq. (4.1) indeed transforms as a scalar
$\mathbb{D}^{\prime ab}\delta{\mathcal{D}}^{\prime}_{ab}+\tilde{\mathcal{E}}^{\prime}_{c}{}^{ab}\delta\Pi^{\prime c}{}_{ab}=\mathbb{D}^{ab}\delta{\mathcal{D}}_{ab}+\tilde{\mathcal{E}}_{c}{}^{ab}\delta\Pi^{c}{}_{ab}$ (4.16)
where each term individually is not coordinate invariant, but their sum is coordinate invariant. These nuances between the different formulations, in particular the tensorial and non-tensorial nature of the different fields and variations, may have implications as to certain features within a quantum theory. In particular, these features may prove vital in developing a Faddeev-Popov-type of gauge fixing procedure.
4.2 Summary of Field Equations for Each Formulation
We see that regardless of which set of fields with respect to we vary the action, we get the same field equations on-shell classically. That is, with one set of fields, the classical equations of motion are linear combinations of the fields for the other set. The equations of motion derived from variation of the projectively invariant and tensorial fields ${\bar{{\mathcal{P}}}}_{ab}$, ${\tilde{C}}^{c}{}_{ab}$, and $g^{mn}$ are
$\mathbb{D}^{ab}=$ $-\frac{d-1}{4J_{0}c\kappa_{0}}g^{ab}-(d-1)g^{ab}\mathcal{K}+2(2d-3)\mathcal{K}^{ab}+\lambda_{0}^{2}\bar{\nabla}_{c}K^{(ab)c}=0$ (4.17)
$\tilde{\mathbb{E}}_{c}{}^{ab}\equiv$ $\mathbb{E}_{c}{}^{ab}-\tfrac{1}{{\rm d}+1}\delta_{c}{}^{(a}\mathbb{E}_{m}{}^{b)m}=0$ (4.18)
$\mathbb{H}_{mn}=$ $\mathcal{H}_{mn}+\tfrac{2\kappa_{0}J_{0}}{{\rm d}+1}g_{mn}{\bar{\nabla}}_{a}{\bar{\nabla}}_{b}\mathbb{D}^{ab}$ (4.19)
$+2\kappa_{0}J_{0}\left[{\bar{\nabla}}_{a}(\tilde{\mathbb{E}}_{(mn)}{}^{a}-\tilde{\mathbb{E}}^{a}{}_{mn})+{\tilde{C}}^{p}{}_{a(m}\tilde{\mathbb{E}}_{|p|n)}{}^{a}+{\tilde{C}}^{p}{}_{a(m}\tilde{\mathbb{E}}_{n)p}{}^{a}-{\tilde{C}}^{p}{}_{a(m}\tilde{\mathbb{E}}^{a}{}_{n)p}\right]$ (4.20)
$=$ 00 0 (4.21)
where
$\mathbb{E}_{c}{}^{ab}=$${\bar{\nabla}}_{c}\left[g^{ab}\left({\mathcal{K}}+\tfrac{1}{4J_{0}\kappa_{0}}\right)-4{\mathcal{K}}^{ab}\right]-{\bar{\nabla}}_{d}\left(g_{cm}{\mathcal{K}}^{m(ab)d}\right)-\lambda_{0}^{2}{\bar{{\mathcal{P}}}}_{dc}K^{(ab)d}$ (4.22)
$\mathcal{H}_{mn}=$ $-\mathcal{K}_{mn}-\kappa_{0}g_{mn}\left[\tfrac{2J_{0}\lambda_{0}^{2}}{{\rm d}+1}{\bar{\nabla}}_{a}\left({\mathcal{K}}^{a}{}_{bcd}K^{bcd}\right)+\mathcal{L}\right]$ (4.23)
$+2\kappa_{0}J_{0}\lambda_{0}^{2}(K_{mcd}K_{n}{}^{cd}+2K_{bcm}K^{bc}{}_{n})+2\kappa_{0}J_{0}(8{\mathcal{K}}_{mb}{\mathcal{K}}^{b}{}_{n}-2{\mathcal{K}}{\mathcal{K}}_{mn})$ (4.24)
$+2\kappa_{0}J_{0}({\mathcal{K}}_{mbcd}{\mathcal{K}}_{n}{}^{bcd}-{\mathcal{K}}^{a}{}_{mcd}{\mathcal{K}}_{an}{}^{cd}-2{\mathcal{K}}^{abc}{}_{m}{\mathcal{K}}_{abcn})$ (4.25)
$\mathcal{L}=$ $-\tfrac{1}{2\kappa_{0}}{\mathcal{K}}+J_{0}\lambda_{0}^{2}K_{bcd}K^{bcd}-J_{0}(\mathcal{K}_{abcd}\mathcal{K}^{abcd}-4\mathcal{K}_{bd}\mathcal{K}^{bd}+\mathcal{K}^{2})~{}~{}~{}.$ (4.26)
The equations of motion derived from variation of the original fields ${\mathcal{D}}_{ab}$, $\Pi^{c}{}_{ab}$, and $g^{ab}$ are
$\mathbb{D}^{ab}=$ $0\displaystyle 0~{}~{}~{},~{}~{}~{}\mathcal{H}_{mn}=$ 0 (4.27)
~$\tilde{\mathcal{E}}_{c}{}^{ab}=$ $\tilde{\mathbb{E}}_{c}{}^{ab}+\left(g_{c}\mathbb{D}^{ab}-\tfrac{1}{{\rm d}+1}\delta_{c}{}^{(a}g_{m}\mathbb{D}^{b)m}\right)=0~{}~{}~{}.$ (4.28)
On shell, these two sets of equations simultaneously reduce to
[…]
Finally, it is useful to express the equations of motion for $g^{mn}$ in the form of Einstein equations. On-shell, this becomes for either formulation
$\hat{R}_{mn}-\tfrac{1}{2}g_{mn}\hat{R}=\kappa_{0}\hat{T}_{mn}$ (4.30)
$\hat{T}_{mn}=-\frac{1}{\kappa_{0}}\left[{\bar{\nabla}}_{a}{\tilde{C}}^{a}{}_{mn}+{\tilde{C}}^{c}{}_{am}{\tilde{C}}^{a}{}_{nc}+({\rm d}-1){\bar{{\mathcal{P}}}}_{mn}\right.$ (4.31)
$\qquad\hskip 10.0pt\left.-\tfrac{1}{2}g_{mn}\left(g^{ab}{\bar{\nabla}}_{c}{\tilde{C}}^{c}{}_{ab}+{\tilde{C}}^{ca}{}_{b}{\tilde{C}}^{b}{}_{ca}+({\rm d}-1){\bar{{\mathcal{P}}}}\right)\right]$
$\qquad\hskip 10.0pt+J_{0}g_{mn}\left[-\tfrac{2\lambda_{0}^{2}}{{\rm d}+1}{\bar{\nabla}}_{a}(K_{bcd}\mathcal{K}^{abcd})-\lambda_{0}^{2}K^{bcd}K_{bcd}+\mathcal{K}_{abcd}\mathcal{K}^{abcd}-4\mathcal{K}^{ab}\mathcal{K}_{ab}+\mathcal{K}^{2}\right]$ (4.32)
$\qquad\hskip 10.0pt+2J_{0}\lambda_{0}^{2}(K_{mcd}K_{n}{}^{cd}+2K_{bcm}K^{bc}{}_{n})+2J_{0}(8{\mathcal{K}}_{mb}{\mathcal{K}}^{b}{}_{n}-2{\mathcal{K}}{\mathcal{K}}_{mn})$
$\qquad\hskip 10.0pt+2J_{0}({\mathcal{K}}_{mbcd}{\mathcal{K}}_{n}{}^{bcd}-{\mathcal{K}}^{a}{}_{mcd}{\mathcal{K}}_{an}{}^{cd}-2{\mathcal{K}}^{abc}{}_{m}{\mathcal{K}}_{abcn})~{}~{}~{}.$
In casting the on-shell equations of motion for $g^{mn}$ in this form, starting with either Eq. (2.71) or Eq. (4.19), the following identities prove useful
$Q_{mn}=$${\bar{\nabla}}_{a}{\tilde{C}}^{a}{}_{mn}+{\tilde{C}}^{c}{}_{am}{\tilde{C}}^{a}{}_{nc}+{\bar{\nabla}}_{[m}\text{\textcent}_{n]}-({\rm d}-1){\bar{\nabla}}_{n}\text{\textcent}_{m}+({\rm d}-1)\text{\textcent}_{m}\text{\textcent}_{n}$ (4.33)
${\mathcal{P}}_{mn}=$ ${\bar{{\mathcal{P}}}}_{mn}+{\bar{\nabla}}_{m}\text{\textcent}_{n}-\text{\textcent}_{m}\text{\textcent}_{n}$ (4.34)
${\mathcal{K}}_{mn}=$ $\hat{R}_{mn}+\tfrac{1}{2}Q_{(mn)}+\tfrac{{\rm d}-1}{2}{\bar{\nabla}}_{(m}\text{\textcent}_{n)}-({\rm d}-1)\text{\textcent}_{m}\text{\textcent}_{n}+({\rm d}-1){\bar{{\mathcal{P}}}}_{mn}$ (4.35)
$=$ $\hat{R}_{mn}+{\bar{\nabla}}_{a}{\tilde{C}}^{a}{}_{mn}+{\tilde{C}}^{c}{}_{am}{\tilde{C}}^{a}{}_{nc}+({\rm d}-1){\bar{{\mathcal{P}}}}_{mn}~{}~{}~{}.$ (4.36)
We note that in the form above the Einstein Equations for TW gravity are manifestly covariant and projectively invariant. Clearly this is true for the stress energy tensor $\hat{T}_{mn}$, though it may seem at first glance that projective invariance is violated on the left hand side. This is not so, indeed the left-hand side is projectively invariant: the Levi-Civita Ricci tensor $\hat{R}_{mn}$ and Levi-Civita Ricci scalar $\hat{R}$ know not of projective transformations as they depend only on the metric and its derivatives. The projective transformation associated with $R_{ab}$ has been fully hidden in the projectively invariant Palatini field ${\tilde{C}}^{c}{}_{ab}$, leaving $\hat{R}_{mn}$ and $\hat{R}$ projectively invariant. Thus the Einstein equations in the above form are indeed covariant and manifestly projectively invariant.
5 Conclusion
The TW action, built out of the projective Einstein-Hilbert and projective Gauss-Bonnet actions, ties projective geometry to string theory and gravity. The projective connection and the metric are independent in the spirit of the Palatini formalism, giving rise to non-metricity degrees of freedom. The TW field equations then have degrees of freedom from the metric, the diffeomorphism sector, and the Palatini field, which may be encapsulated in different choices of combinations of fields that we have shown are equivalent up to different linear combinations. In this way projective geometry can influence General Relativity by providing sources in the energy-momentum tensor, giving a possible geometric origin to dark energy, dark matter, inflation, and other physical phenomenon.
This paper analyzed pure TW gravity, with no matter or other energy sources: all of the aforementioned applications are of TW gravity itself. Interactions with matter and energy, from standard model sources for instance, are straightforward to include: source terms would appear on the the right hand sides of Eqs. (4.29) rather than being equal to zero. In the form of Einsteins Equations (4.30), interactions with other sources would appear as additions to the stress-energy tensor in the usual way. Indeed, this was already discussed in the previous paper [20]. We look forward to investigating the form of such interactions, in this covariant and projectively invariant formulation, in a future work.
As the equations of motion are different off-shell as derived with respect to different sets of fields, this will have interesting implications at the quantum level in future works. If for instance different gauge fixing procedures produce different results, this could perhaps be related to the trace anomaly in string theory. To iron out such details, the development of a Fadeev-Popov-type of gauge fixing procedure for TW gravity is under consideration. TW gravity itself may be an effective action of a deeper, underlying theory like string theory. Logarithmic running of the coupling constants is already present in the classical theory as demonstrated in this and previous works. Now having a formulation that is both tensorial and projectively invariant will provide a platform for these further investigations.
6 Acknowledgments
The authors would like to thank the Diffeomorphisms and Geometry Research Group and the High Energy Physics Group at the University of Iowa for constructive conversations throughout this work. In particular, V.G.J. Rodgers was the source of much inspiration and elucidation.
Appendix A Conventions and Definitions
Our index convention is as in [20]: Latin indices $a,b,c,\dots$ are over the ${\rm d}$-dimensional manifold $\mathcal{M}$ and Greek indices $\alpha,\beta,\gamma,\dots$ excluding $\lambda$ are over the ${\rm d}+1$-dimensional manifold $\mathcal{N}$, also called the Thomas cone. The connection and curvature tensors on the Thomas cone are succinctly listed in Sections 2.1 and 2.2. The various objects on the Manifold $\mathcal{M}$ are located throughout the paper, so we succinctly list them all here. Parentheses and brackets denote symmetrization and antisymmetrization, respectively. For example, for some arbitrary rank 2 tensor $T$ we define
$T_{(ab)}=T_{ab}+T_{ba}~{}~{}~{},~{}~{}~{}T_{[ab]}=T_{ab}-T_{ba}$ (A.1)
where we do not include a factor of $\frac{1}{2}$.
The different connections used throughout are related as
$\bar{\Gamma}^{a}{}_{bc}$ $=\Gamma^{a}{}_{bc}-\delta^{a}{}_{(b}\text{\textcent}_{c)}=\Pi^{a}_{\ bc}-\delta^{a}_{\ (b}g_{c)}~{}~{}~{}.$ (A.2)
The piecemeal parts used to construct fields are
$\alpha_{m}=$ $-\tfrac{1}{d+1}\Gamma^{e}_{\ em}$ (A.3)
$\alpha_{\mu}=$ $(\alpha_{m},\lambda^{-1})$ (A.4)
$g_{m}=$ $-\tfrac{1}{d+1}\partial_{m}\log\sqrt{|g|}=-\tfrac{1}{d+1}\hat{\Gamma}^{e}{}_{me}$ (A.5)
$g_{\mu}=$ $(g_{m},\lambda^{-1})$ (A.6)
¢$\text{\textcent}_{m}=$ $\ g_{m}-\alpha_{m}$ (A.7)
with $|g|=\det g_{ab}$ where $g_{ab}$ is the metric on the manifold $\mathcal{M}$. Each connection can be expanded in terms of the Levi-Civita connection $\hat{\Gamma}^{a}{}_{bc}$ and Palatini tensor $C^{a}{}_{bc}$ as
$\bar{\Gamma}^{a}{}_{bc}=$ $\hat{\Gamma}^{a}{}_{bc}+\tilde{C}^{a}{}_{bc}$ (A.8)
$\Gamma^{a}{}_{bc}=$ $\hat{\Gamma}^{a}{}_{bc}+C^{a}{}_{bc}$ (A.9)
$\Pi^{a}{}_{bc}=$ $\hat{\Gamma}^{a}{}_{bc}+\delta^{a}{}_{(b}g_{c)}+\tilde{C}^{a}{}_{bc}~{}~{}~{}.$ (A.10)
All connections as well as the Palatini tensor are symmetric in their lower two indices as for instance
$C^{a}{}_{bc}=C^{a}{}_{cb}~{}~{}~{}.$ (A.11)
The Levi-Civita connection’s components are the canonical Christoffel symbols and $\tilde{C}^{a}{}_{bc}$ is the traceless part of the arbitrary Palatini tensor
$\hat{\Gamma}^{c}{}_{ab}=$ $\frac{1}{2}g^{cm}\left(\partial_{(a}g_{b)m}-\partial_{m}g_{ab}\right)$ (A.12)
$\tilde{C}^{a}{}_{bc}=$ $C^{a}{}_{bc}-\tfrac{1}{{\rm d}+1}\delta^{a}{}_{(b}C^{m}{}_{c)m}=C^{a}{}_{bc}-\delta^{a}{}_{(b}\text{\textcent}_{c)}~{}~{}~{}.$ (A.13)
In matrix form, the metric and its inverse on the Thomas cone are
[…]
More succinctly, we have
$G_{\alpha\beta}$ $=\delta^{a}_{\,\,\alpha}\delta^{b}_{\,\,\beta}\,g_{ab}-\lambda_{0}^{2}g_{\alpha}g_{\beta}$ (A.20)
$G^{\alpha\beta}$ $=g^{ab}(\delta^{\alpha}_{\ a}-g_{a}\Upsilon^{\alpha})(\delta^{\beta}_{\ b}-g_{b}\Upsilon^{\alpha})-\lambda_{0}^{-2}\Upsilon^{\alpha}\Upsilon^{\beta}$ (A.21)
$\Upsilon^{\alpha}=$ $(0,0,\dots,0,\lambda)~{}~{}~{}.$ (A.22)
On the Manifold $\mathcal{M}$, the curvature tensors are
$\hat{R}^{a}{}_{bcd}=$ $\hat{\Gamma}^{a}{}_{b[d,c]}+\hat{\Gamma}^{m}{}_{b[d}\hat{\Gamma}^{a}{}_{c]m}$ (A.23)
$Q^{a}_{\ bcd}=$$\hat{\nabla}_{c}C^{a}_{\ bd}-\hat{\nabla}_{d}C^{a}_{\ bc}+C^{a}_{\ ce}C^{e}_{\ bd}-C^{a}_{\ de}C^{e}_{\ bc}$ (A.24)
$\bar{R}^{a}{}_{bcd}=$ $\bar{\Gamma}^{a}{}_{b[d,c]}+\bar{\Gamma}^{m}{}_{b[d}\bar{\Gamma}^{a}{}_{c]m}$ (A.25)
$R^{a}{}_{bcd}=$ $\Gamma^{a}{}_{b[d,c]}+\Gamma^{m}{}_{b[d}\Gamma^{a}{}_{c]m}$ (A.26)
$\qquad$== $\hat{R}^{a}_{\ bcd}+Q^{a}_{\ bcd}$ (A.27)
$\mathcal{R}^{a}{}_{bcd}=$ $\Pi^{a}{}_{b[d,c]}+\Pi^{m}{}_{b[d}\Pi^{a}{}_{c]m}$ (A.28)
$\mathcal{K}^{a}{}_{bcd}=$ $R^{a}{}_{bcd}+\delta^{a}_{\ [\,c}{}\mathcal{P}_{d]b}-\delta^{a}{}_{b}\mathcal{P}_{[cd]}$ (A.29)
$\qquad$== $\mathcal{R}^{a}_{\ bcd}+\delta^{a}_{\ [c}\mathcal{D}_{d]b}$ (A.30)
$=$ $\bar{R}^{a}{}_{bcd}+\delta^{a}{}_{[c}\bar{\mathcal{P}}_{d]b}~{}~{}~{}.$ (A.31)
The projective Cotton-York tensor is
$\mathcal{K}_{bcd}=$$\breve{\nabla}_{[c}\mathcal{D}_{d]b}+g_{a}\mathcal{K}^{a}{}_{bcd}$
$=$ $\nabla_{[c}\mathcal{P}_{d]b}+\text{\textcent}_{a}\mathcal{K}^{a}{}_{bcd}$
$=$ ¯$\bar{\nabla}_{[c}\bar{\mathcal{P}}_{d]b}~{}~{}~{}.$ (A.32)
Ricci tensors are always constructed by contracting over the first and third indices, and Ricci scalars are the trace of their associated Ricci tensors as seen below
$\mathcal{K}_{bd}=$ $\mathcal{K}^{a}_{\ bad}=\mathcal{R}_{bd}+(d-1)\mathcal{D}_{bd}=R_{bd}+d\mathcal{P}_{db}-\mathcal{P}_{bd}$
$=$ $\bar{R}_{mn}+(d-1)\bar{\mathcal{P}}_{mn}$ (A.33)
$\mathcal{K}=$ $g^{ab}\mathcal{K}_{ab}=\mathcal{R}+(d-1)\mathcal{D}=R+(d-1)\mathcal{P}$
$=$ $\bar{R}+(d-1)\bar{\mathcal{P}}$ (A.34)
$\mathcal{R}_{bd}=$ $\mathcal{R}^{a}_{\ bad}$ (A.35)
$R_{bd}=$ $R^{a}_{\ bad}=\hat{R}^{a}_{\ bad}+Q^{a}_{\ bad}$
$\qquad$== $\hat{R}_{bd}+Q_{bd}$ (A.36)
$\bar{R}_{bd}=$ $\bar{R}^{a}{}_{bad}$
$\qquad$== $\tfrac{1}{2}R_{(bd)}+(d-1)\nabla_{(b}\text{\textcent}_{d)}+(d-1)\text{\textcent}_{b}\text{\textcent}_{d}$ (A.37)
$\mathcal{R}=$ $g^{bd}\mathcal{R}_{bd}$ (A.38)
$R=$ $g^{bd}R_{bd}=g^{bd}\hat{R}_{bd}+g^{bd}Q_{bd}=\hat{R}+Q$ (A.39)
$\bar{R}=$ $g^{bd}\bar{R}_{bd}=R+2(d-1)\nabla^{m}\text{\textcent}_{m}+(d-1)\text{\textcent}_{m}\text{\textcent}^{m}$ (A.40)
$\mathcal{P}=$ $g^{ab}\mathcal{P}_{ab}~{}~{}~{},~{}~{}~{}\mathcal{D}=g^{ab}\mathcal{D}_{ab}~{}~{}~{},~{}~{}~{}Q=g^{ab}Q_{ab}~{}~{}~{}.$ (A.41)
The Diffeomorphism field and its related tensors are
$\bar{\mathcal{P}}_{mn}=$ $\mathcal{D}_{mn}-\breve{\nabla}_{m}g_{n}-g_{m}g_{n}=\mathcal{P}_{mn}-\nabla_{m}\text{\textcent}_{n}-\text{\textcent}_{m}\text{\textcent}_{n}$ (A.42)
$\bar{\mathcal{P}}=g^{mn}\bar{\mathcal{P}}_{mn}=$ $\mathcal{D}--g^{mn}\breve{\nabla}_{m}g_{n}-g^{mn}g_{m}g_{n}=\mathcal{P}-\nabla^{m}\text{\textcent}_{m}-\text{\textcent}_{m}\text{\textcent}^{m}~{}~{}~{}.$ (A.43)
The following are used to write the equations of motion of Thomas-Whitehead gravity in their original form as in  [20]
$\hat{\mathcal{K}}_{d}^{\ abc}=$ $\mathcal{K}^{\bar{d}}_{\ \bar{a}\bar{b}\bar{c}}\mathcal{G}_{d\bar{d}}^{\ \ \ a\bar{a}[b|\bar{b}|c]\bar{c}}$ (A.44)
$\mathcal{G}_{d\bar{d}}^{\ \ \ a\bar{a}b\bar{b}c\bar{c}}=$ $G_{d\bar{d}}g^{a\bar{a}}g^{b\bar{b}}g^{c\bar{c}}-4\delta^{b}_{\ d}\delta^{\bar{b}}_{\ \bar{d}}g^{a\bar{a}}g^{c\bar{c}}+\delta^{b}_{\ d}\delta^{\bar{b}}_{\ \bar{d}}g^{ac}g^{\bar{a}\bar{c}}$ (A.45)
$\breve{\mathcal{K}}^{abc}=$ $g^{ap}g^{bm}g^{cn}\breve{\nabla}_{[m}{\mathcal{D}}_{n]p}~{}~{}~{}.$ (A.46)
Appendix B Recasting the Original TW Field Equations in a Manifestly Tensorial Form
To rewrite the Field equations in terms of $g_{a}$ instead of explicit factors of $\sqrt{|g|}$, the following property will be useful
$\tfrac{1}{\sqrt{|g|}}\breve{\nabla}_{a}(\sqrt{|g|}Anything)=$ (˘$(\breve{\nabla}_{a}-({\rm d}+1)g_{a})Anything~{}~{}~{}.$ (B.1)
B.1 Recasting $\mathbb{D}^{ab}$ in a Manifestly Tensorial Form
Equation (B.1) allows us to expand terms such as
$\tfrac{1}{\sqrt{|g|}}\breve{\nabla}_{a}(\sqrt{|g|}K^{(ab)c})=$ (˘$(\breve{\nabla}_{a}-({\rm d}+1)g_{a})K^{(ab)c}$ (B.2)
$=$ $(\nabla_{a}-({\rm d}+2)\text{\textcent}_{c})K^{(ab)c}+g_{c}K^{(ab)c}$ (B.3)
where we have used Eqs. (A.2) and (A.7) to produce the last line. It will also be of use to expand Eqs. (A.44) and (A.46)
$\hat{{\mathcal{K}}}_{c}{}^{(ab)d}=$ $2{\mathcal{K}}_{c}{}^{(ab)d}-2\lambda_{0}^{2}g_{c}g_{m}{\mathcal{K}}^{m(ab)d}+\delta^{d}{}_{c}(8{\mathcal{K}}^{ab}-2g^{ab}{\mathcal{K}})+\delta^{(a}{}_{c}(g^{b)d}{\mathcal{K}}-4{\mathcal{K}}^{b)d})$ (B.4)
$\breve{{\mathcal{K}}}^{(ab)c}=$ $K^{(ab)c}-g_{m}{\mathcal{K}}^{m(ab)c}~{}~{}~{}.$ (B.5)
In expanding $\breve{{\mathcal{K}}}^{(ab)c}$, we have made use of the definition of the projective Cotton-York tensor from Eq. (2.40). Contracting indices in Eq. B.4 and simplifying yields the following useful result
$\tfrac{1}{2}\hat{{\mathcal{K}}}_{c}{}^{(ab)c}=-\lambda_{0}^{2}g_{c}g_{m}{\mathcal{K}}^{m(ab)c}+2(2{\rm d}-3){\mathcal{K}}^{ab}-({\rm d}-1)g^{ab}{\mathcal{K}}~{}~{}~{}.$ (B.6)
Recall the field equation $\mathbb{D}^{ab}$ in its manifestly projectively invariant form
$\mathbb{D}^{ab}\equiv-\tfrac{({\rm d}-1)}{4\kappa_{0}J_{0}}g^{ab}+\tfrac{1}{2}\hat{\mathcal{K}}_{c}^{\ (ab)c}+\tfrac{\lambda_{0}^{2}}{\sqrt{|g|}}\breve{\nabla}_{c}(\sqrt{|g|}K^{(ab)c})-\lambda_{0}^{2}g_{c}\breve{\mathcal{K}}^{(ab)c}=0~{}~{}~{}.$ (2.59)
Utilizing the expansions in Eqs. (B.2), (B.6), and (B.5) and simplifying casts Eq. $\mathbb{D}^{ab}$ into its manifestly tensorial form
$\mathbb{D}^{ab}=$ $-\tfrac{d-1}{4J_{0}\kappa_{0}}g^{ab}-(d-1)g^{ab}\mathcal{K}+2(2d-3)\mathcal{K}^{ab}+\lambda_{0}^{2}(\nabla_{c}-({\rm d}+2)\text{\textcent}_{c})K^{(ab)c}=0.$ (2.70)
Utilizing Eq. (3.4), it is straightforward to show the last term collapses to an overall ¯$\bar{\nabla}_{c}$ derivative. This reduces the equation to the covariant and manifestly projectively invariant form
$\mathbb{D}^{ab}=$ $-\tfrac{d-1}{4J_{0}\kappa_{0}}g^{ab}-(d-1)g^{ab}\mathcal{K}+2(2d-3)\mathcal{K}^{ab}+\lambda_{0}^{2}\bar{\nabla}_{c}K^{(ab)c}~{}~{}~{}.$ (3.40)
B.2 Recasting $\mathbb{H}^{ab}$ in a Manifestly Tensorial Form
Using Eqs. (B.1), (A.2), and (A.7) we find the following identity that we will find useful in this section
$\tfrac{1}{\sqrt{|g|}}\breve{\nabla}_{a}(\sqrt{|g|}{\mathcal{K}}^{a}{}_{bcd}K^{bcd})=$ $(\nabla_{a}-({\rm d}+1)\text{\textcent}_{a}){\mathcal{K}}^{a}{}_{bcd}K^{bcd}~{}~{}~{}.$ (B.7)
Beginning with the equation of motion $\mathcal{H}_{mn}$
$\mathcal{H}_{mn}=$ $-\mathcal{K}_{mn}-\kappa_{0}g_{mn}\left[\tfrac{2J_{0}\lambda_{0}^{2}}{({\rm d}+1)\sqrt{|g|}}\breve{\nabla}_{a}(\sqrt{|g|}{\mathcal{K}}^{a}{}_{bcd}K^{bcd})+\mathcal{L}\right]$ (B.8)
$+2\kappa_{0}J_{0}\lambda_{0}^{2}(K_{mcd}K_{n}{}^{cd}+2K_{bcm}K^{bc}{}_{n})+2\kappa_{0}J_{0}(8{\mathcal{K}}_{mb}{\mathcal{K}}^{b}{}_{n}-2{\mathcal{K}}{\mathcal{K}}_{mn})$ (B.9)
$+2\kappa_{0}J_{0}({\mathcal{K}}_{mbcd}{\mathcal{K}}_{n}{}^{bcd}-{\mathcal{K}}^{a}{}_{mcd}{\mathcal{K}}_{an}{}^{cd}-2{\mathcal{K}}^{abc}{}_{m}{\mathcal{K}}_{abcn})$ (B.10)
[…]
$\mathcal{L}=$ $-\tfrac{1}{2\kappa_{0}}{\mathcal{K}}+J_{0}\lambda_{0}^{2}K_{bcd}K^{bcd}-J_{0}(\mathcal{K}_{abcd}\mathcal{K}^{abcd}-4\mathcal{K}_{bd}\mathcal{K}^{bd}+\mathcal{K}^{2})$ (2.69)
we use Eq. (B.7) along with Eqs. (A), (A), (A), and (A.39) and simplify to cast $\mathcal{H}_{mn}$ into the following form
$\hat{R}_{mn}-\tfrac{1}{2}g_{mn}\hat{R}=\kappa_{0}\hat{T}_{mn}$ (B.11)
$\hat{T}_{mn}=-\frac{1}{\kappa_{0}}\Big{(}Q_{mn}-\tfrac{1}{2}g_{mn}Q+({\rm d}-1)\mathcal{P}_{mn}-{\rm d}{\mathcal{P}}_{[mn]}-\tfrac{{\rm d}-1}{2}g_{mn}\mathcal{P}\Big{)}$
$\qquad-J_{0}g_{mn}\left[\lambda_{0}^{2}K_{bcd}K^{bcd}-({\mathcal{K}}_{abcd}{\mathcal{K}}^{abcd}-4{\mathcal{K}}_{ab}{\mathcal{K}}^{ab}+{\mathcal{K}}^{2})\right]$ (B.12)
$\qquad+g_{mn}2J_{0}\lambda_{0}^{2}(-\tfrac{1}{{\rm d}+1}\nabla_{a}+\text{\textcent}_{a})(K_{bcd}\mathcal{K}^{abcd})$ (B.13)
$\qquad+2J_{0}\lambda_{0}^{2}(K_{mcd}K_{n}{}^{cd}+2K_{bcm}K^{bc}{}_{n})+2J_{0}(8{\mathcal{K}}_{mb}{\mathcal{K}}^{b}{}_{n}-2{\mathcal{K}}{\mathcal{K}}_{mn})$ (B.14)
$\qquad+2J_{0}({\mathcal{K}}_{mbcd}{\mathcal{K}}_{n}{}^{bcd}-{\mathcal{K}}^{a}{}_{mcd}{\mathcal{K}}_{an}{}^{cd}-2{\mathcal{K}}^{abc}{}_{m}{\mathcal{K}}_{abcn})]~{}~{}~{}.$
Combining lines one and two of $\hat{T}_{mn}$ and simplifying results in the covariant form of Einstein’s equations
$\hat{R}_{mn}-\tfrac{1}{2}g_{mn}\hat{R}=\kappa_{0}\hat{T}_{mn}$ (2.71)
$\hat{T}_{mn}=-\frac{1}{\kappa_{0}}\Big{(}Q_{mn}-\tfrac{1}{2}g_{mn}Q+({\rm d}-1)\mathcal{P}_{mn}-{\rm d}{\mathcal{P}}_{[mn]}-\tfrac{{\rm d}-1}{2}g_{mn}\mathcal{P}\Big{)}$
$\qquad+J_{0}g_{mn}\left[2\lambda_{0}^{2}(-\tfrac{1}{{\rm d}+1}\nabla_{a}+\text{\textcent}_{a})K_{bcd}\mathcal{K}^{abcd}-\lambda_{0}^{2}K^{bcd}K_{bcd}\right.$ (B.15)
$\left.\qquad\qquad\qquad+\mathcal{K}_{abcd}\mathcal{K}^{abcd}-4\mathcal{K}^{ab}\mathcal{K}_{ab}+\mathcal{K}^{2}\right]+2J_{0}\lambda_{0}^{2}(K_{mcd}K_{n}{}^{cd}+2K_{bcm}K^{bc}{}_{n})$
$\qquad+2J_{0}(8{\mathcal{K}}_{mb}{\mathcal{K}}^{b}{}_{n}-2{\mathcal{K}}{\mathcal{K}}_{mn})+2J_{0}({\mathcal{K}}_{mbcd}{\mathcal{K}}_{n}{}^{bcd}-{\mathcal{K}}^{a}{}_{mcd}{\mathcal{K}}_{an}{}^{cd}-2{\mathcal{K}}^{abc}{}_{m}{\mathcal{K}}_{abcn}).$
B.3 Recasting $\tilde{E}_{c}{}^{ab}$ in terms of Tensorial Objects to Demonstrate it is not a Tensor
Using Eqs. (B.1) and (B.2) it is straightforward to prove the following identities that we will find useful in this section
$\tfrac{1}{\sqrt{|g|}}\breve{\nabla}_{c}(\sqrt{|g|}T^{ab})=$ $(\nabla_{c}-2\text{\textcent}_{c})T^{ab}-\delta^{(a}{}_{c}\text{\textcent}_{m}T^{b)m}-({\rm d}-1)g_{c}T^{ab}+\delta^{(a}{}_{c}g_{m}T^{b)m}$ (B.16)
$\tfrac{1}{\sqrt{|g|}}\breve{\nabla}_{d}(\sqrt{|g|}K_{c}{}^{(ab)d})=$ $(\nabla_{d}-({\rm d}+1)\text{\textcent}_{d}){\mathcal{K}}_{c}{}^{(ab)d}-2\text{\textcent}_{c}{\mathcal{K}}^{ab}+2g_{c}{\mathcal{K}}^{ab}$ (B.17)
$\tfrac{1}{\sqrt{|g|}}\breve{\nabla}_{d}(\sqrt{|g|}g_{c}K^{(ab)d})=$ $K^{(ab)d}(\breve{\nabla}_{c}g_{d}+g_{c}g_{d})+g_{c}(\nabla_{d}-({\rm d}+2)\text{\textcent}_{d})K^{(ab)d}~{}~{}~{}.$ (B.18)
In the above, $T^{ab}=T^{ba}$ represents an arbitrary symmetric tensor. Starting with the original form of the equations of motion
~$\tilde{\mathcal{E}}_{c}{}^{ab}\equiv\mathcal{E}_{c}^{\ ab}-\tfrac{1}{d+1}\delta_{c}^{\ (a}\mathcal{E}_{d}^{\ b)d}=0$ (2.63)
$\mathcal{E}_{c}^{\ ab}=\mathcal{E}_{c}^{\ ba}\equiv\tfrac{1}{4\kappa_{0}J_{0}\sqrt{|g|}}\breve{\nabla}_{c}(\sqrt{|g|}g^{ab})-\tfrac{1}{2\sqrt{|g|}}\breve{\nabla}_{d}(\sqrt{|g|}\hat{\mathcal{K}}_{c}^{\ (ab)d})\quad$
$\qquad\qquad\qquad\qquad+\tfrac{\lambda_{0}^{2}}{\sqrt{|g|}}\breve{\nabla}_{d}(\sqrt{|g|}g_{c}\breve{\mathcal{K}}^{(ab)d})-\lambda_{0}^{2}K^{(ab)d}\mathcal{D}_{dc}~{}~{}~{}$ (2.4)
we use Eqs. (B.4) and (B.5) to substitute for $\hat{{\mathcal{K}}}_{c}{}^{(ab)d}$ and $\breve{{\mathcal{K}}}^{(ab)d}$. After some simplification this results in
$\mathcal{E}_{c}^{\ ab}=$ $\tfrac{1}{\sqrt{|g|}}\breve{\nabla}_{c}\left[\sqrt{|g|}\left((\tfrac{1}{4J_{0}\kappa_{0}}+\mathcal{K})g^{ab}-4\mathcal{K}^{ab}\right)\right]-\tfrac{1}{\sqrt{|g|}}\breve{\nabla}_{d}(\sqrt{|g|}\mathcal{K}_{c}^{\ (ab)d})\quad$ (B.19)
$+\lambda_{0}^{2}\left[\tfrac{1}{\sqrt{|g|}}\breve{\nabla}_{d}(\sqrt{|g|}g_{c}K^{(ab)d})-{\mathcal{D}}_{cd}K^{(ab)d}\right]$ (B.20)
$-\tfrac{1}{2}\delta^{(a}{}_{c}\tfrac{1}{\sqrt{|g|}}\breve{\nabla}_{d}\left[\sqrt{|g|}\left(g^{b)d}{\mathcal{K}}-4{\mathcal{K}}^{b)d}\right)\right]~{}~{}~{}.$ (B.21)
Next, we substitute using Eqs. (B.16), (B.17), and (B.18) and simplify to the final form
$\mathcal{E}_{c}{}^{ab}=$ $\mathbb{E}_{c}{}^{ab}+g_{c}\mathbb{D}^{ab}+\delta^{(a}{}_{c}\mathfrak{g}^{b)}$ (2.73)
$\mathbb{E}_{c}{}^{ab}\equiv$ $(\nabla_{c}-2\text{\textcent}_{c})\mathscr{K}^{ab}-\delta^{(a}{}_{c}\text{\textcent}_{d}\mathscr{K}^{b)d}-[(\nabla_{d}-({\rm d}+1)\text{\textcent}_{d})\mathcal{K}_{c}{}^{(ab)d}-2\text{\textcent}_{c}\mathcal{K}^{ab}]$ (B.22)
$-\lambda_{0}^{2}K^{(ab)d}(\mathcal{P}_{cd}-\nabla_{c}\text{\textcent}_{d}-\text{\textcent}_{c}\text{\textcent}_{d})$ (2.74)
$\mathscr{K}^{ab}\equiv$ $(\tfrac{1}{4J_{0}\kappa_{0}}+\mathcal{K})g^{ab}-4\mathcal{K}^{ab}$ (2.76)
$\mathfrak{g}^{b}\equiv$ $\tfrac{1}{4J_{0}\kappa_{0}}g_{d}g^{bd}-\tfrac{1}{2}(\nabla_{d}-({\rm d}+3)\text{\textcent}_{d})(g^{bd}\mathcal{K}-4\mathcal{K}^{bd})$ (2.77)
where in the last line of the definition of $\mathbb{E}_{c}{}^{ab}$, we have made use of Eq. (A.42) to switch between ${\mathcal{D}}_{cd}$ and ${\mathcal{P}}_{cd}$. Subtracting the trace then results in the full equation of motion
~$\tilde{\mathcal{E}}_{c}{}^{ab}=$ $\mathbb{E}_{c}{}^{ab}-\tfrac{1}{{\rm d}+1}\delta^{(a}{}_{c}\mathbb{E}_{m}{}^{b)m}+g_{c}\mathbb{D}^{ab}-\tfrac{1}{{\rm d}+1}\delta^{(a}{}_{c}g_{m}\mathbb{D}^{b)m}=0~{}~{}~{}.$ (2.78)
Appendix C Proofs
C.1 Proof of Eq. (3.63)
We start by calculating the trace $\mathbb{E}_{b}{}^{ab}$ via Eq. (3.50). Inserting this into Eq. (3.63) results in
$V^{a}=$ $\frac{2}{{\rm d}+1}\left\{{\bar{\nabla}}_{b}\left[g^{ab}({\mathcal{K}}+\tfrac{1}{J_{0}\kappa_{0}})-4{\mathcal{K}}^{ab}\right]-{\bar{\nabla}}_{d}(g_{bm}{\mathcal{K}}^{m(ab)d})-\lambda_{0}^{2}{\bar{{\mathcal{P}}}}_{db}K^{(ab)d}\right\}~{}~{}~{}.$ (C.1)
Using the fact that $g_{bm}{\mathcal{K}}^{mbad}$ vanishes trivially from its definition, the second term results in an additional projective Ricci term. Combining like terms and simplifying, we then have
$V^{a}=\tfrac{1}{{\rm d}+1}{\bar{\nabla}}_{b}\left[-({\rm d}-1)g^{ab}(\tfrac{1}{4J_{0}\kappa_{0}}+{\mathcal{K}})+2(2{\rm d}-3){\mathcal{K}}^{ab}\right]-\tfrac{2\lambda_{0}^{2}}{{\rm d}+1}{\bar{{\mathcal{P}}}}_{db}K^{bad}~{}~{}~{}.$ (C.2)
By inspection of Eq. (3.40), we see that the term in brackets is equivalent to $\mathbb{D}^{ab}-\lambda_{0}^{2}{\bar{\nabla}}_{c}K^{(ab)c}$. Substituting this and simplifying results in the second line of Eq. (3.63)
$V^{a}=$ $\tfrac{1}{{\rm d}+1}{\bar{\nabla}}_{b}\mathbb{D}^{ab}-\tfrac{\lambda_{0}^{2}}{{\rm d}+1}\left({\bar{\nabla}}_{b}{\bar{\nabla}}_{c}K^{(ab)c}+2{\bar{{\mathcal{P}}}}_{db}K^{bad}\right)~{}~{}~{}.$ (C.3)
C.2 Proof of Eqs. (3.65) and (3.66)
We expand the variation of the Levi-Civita connection in terms of the metric
$\delta\hat{\Gamma}^{c}{}_{ab}=$ $\tfrac{1}{2}\delta\left[g^{cm}(g_{m(a,b)}-g_{ab,m})\right]$ (C.4)
where a comma denotes a partial derivative. Using Leibniz’s rule for the variation leads to
$\delta\hat{\Gamma}^{c}{}_{ab}=$ $\tfrac{1}{2}\left[\delta g^{cm}(g_{m(a,b)}-g_{ab,m})+g^{cm}(\delta g_{m(a,b)}-\delta g_{ab,m})\right]~{}~{}~{}.$ (C.5)
Switching to Levi-Civita covariant derivatives in the second set of parenthesis yields
$\delta\hat{\Gamma}^{c}{}_{ab}=$ $\tfrac{1}{2}\big{[}\delta g^{cm}(g_{m(a,b)}-g_{ab,m})+g^{cm}(\hat{\nabla}_{(a}\delta g_{b)m}-\hat{\nabla}_{m}\delta g_{ab})$ (C.6)
$+g^{cm}\hat{\Gamma}^{p}{}_{(ab)}\delta g_{pm}+g^{cm}\hat{\Gamma}^{p}{}_{m(a}\delta g_{b)p}-g^{cm}\hat{\Gamma}^{p}{}_{m(a}\delta g_{b)p}\big{]}~{}~{}~{}.$ (C.7)
Cancelling the last two terms, expanding the Levi-Civita connection in terms of the metric in the first term on the second line, and rearranging we have
$\delta\hat{\Gamma}^{c}{}_{ab}=$ $\tfrac{1}{2}g^{cm}(\hat{\nabla}_{(a}\delta g_{b)m}-\hat{\nabla}_{m}\delta g_{ab})$ (C.8)
$+\tfrac{1}{2}\big{[}\delta g^{cm}(g_{m(a,b)}-g_{ab,m})+\delta g_{pm}g^{cm}g^{pn}(g_{n(a,b)}-g_{ab,n})\big{]}$ (C.9)
raising the indices on the variation in the last term, at the cost of a minus sign, yields
$\delta\hat{\Gamma}^{c}{}_{ab}=$ $\tfrac{1}{2}g^{cm}(\hat{\nabla}_{(a}\delta g_{b)m}-\hat{\nabla}_{m}\delta g_{ab})$ (C.10)
$+\tfrac{1}{2}\big{[}\delta g^{cm}(g_{m(a,b)}-g_{ab,m})-\delta g^{cn}(g_{n(a,b)}-g_{ab,n})\big{]}~{}~{}~{}.$ (C.11)
Relabeling indices, the second line vanishes, resulting in
$\delta\hat{\Gamma}^{c}{}_{ab}=$ $\tfrac{1}{2}g^{cm}(\hat{\nabla}_{(a}\delta g_{b)m}-\hat{\nabla}_{m}\delta g_{ab})~{}~{}~{}.$ (C.12)
Contracting indices and simplifying, we find for the variation of the trace
$\delta\hat{\Gamma}^{m}{}_{am}=$ $-\tfrac{1}{2}g_{mn}\hat{\nabla}_{a}\delta g^{mn}~{}~{}~{}.$ (C.13)
Next, we take a Lagrangian variation of the form
$\sqrt{g}A_{c}{}^{ab}\delta\hat{\Gamma}^{c}{}_{ab}$ (C.14)
where $A_{c}{}^{ab}=A_{c}{}^{ba}$ is an arbitrary tensor symmetric in its last two indices. Substituting Eq. (C.12) we have
$\sqrt{g}A_{c}{}^{ab}\delta\hat{\Gamma}^{c}{}_{ab}=$ $\sqrt{g}A_{c}{}^{ab}\tfrac{1}{2}g^{cm}(\hat{\nabla}_{(a}\delta g_{b)m}-\hat{\nabla}_{m}\delta g_{ab})~{}~{}~{}.$ (C.15)
Up to a total derivative, this is equivalent to the following
$\sqrt{g}A_{c}{}^{ab}\delta\hat{\Gamma}^{c}{}_{ab}=$ $\sqrt{g}\tfrac{1}{2}g^{cm}(\delta g_{ab}\hat{\nabla}_{m}A_{c}{}^{ab}-2\delta g_{bm}\hat{\nabla}_{a}A_{c}{}^{ab})~{}~{}~{}.$ (C.16)
Contracting with the metric, relabeling indices, and simplifying, we have
$\sqrt{g}A_{c}{}^{ab}\delta\hat{\Gamma}^{c}{}_{ab}=$ $\sqrt{g}\tfrac{1}{2}\left[\hat{\nabla}_{a}(A^{amn}-2A^{mna})\right]\delta g_{mn}~{}~{}~{}.$ (C.17)
Raising indices on the variation of the metric at the cost of a minus sign and using symmetry of the metric indices to symmetrize indices on $A_{mn}{}^{a}$ yields Eq. (3.65)
$\sqrt{g}A_{c}{}^{ab}\delta\hat{\Gamma}^{c}{}_{ab}=$ $\sqrt{g}\tfrac{1}{2}\left[\hat{\nabla}_{a}(A_{(mn)}{}^{a}-A^{a}{}_{mn})\right]\delta g^{mn}~{}~{}~{}.$ (C.18)
Doing the corresponding calculation for the variation of the trace contracted with an arbitrary vector $A^{a}$ quickly yields Eq. (3.66)
$\sqrt{g}A^{a}\delta\hat{\Gamma}^{m}{}_{am}=$ $\sqrt{g}A^{a}(-\tfrac{1}{2}g_{mn}\hat{\nabla}_{a}\delta g^{mn})$ (C.19)
$=$ $\sqrt{g}\tfrac{1}{2}g_{mn}(\hat{\nabla}_{a}A^{a})\delta g^{mn}$ (C.20)
$=$ $\sqrt{g}\tfrac{1}{2}g_{mn}({\bar{\nabla}}_{a}A^{a})\delta g^{mn}~{}~{}~{}.$ (C.21)
The second equality holds, again, up to a total derivative. In going from the second to third lines, we have used the interchangeability of the divergences of ^$\hat{\nabla}_{a}$ and ¯${\bar{\nabla}}_{a}$, as proved in Eq. (3.22).
C.3 Proof of Eq. (3.70)
Taking the divergence of Eq. (C.3) yields
¯${\bar{\nabla}}_{a}V^{a}=\tfrac{1}{{\rm d}+1}{\bar{\nabla}}_{a}{\bar{\nabla}}_{b}\mathbb{D}^{ab}-\tfrac{\lambda_{0}^{2}}{{\rm d}+1}\left[{\bar{\nabla}}_{a}{\bar{\nabla}}_{b}{\bar{\nabla}}_{c}K^{(ab)c}+2{\bar{\nabla}}_{a}({\bar{{\mathcal{P}}}}_{db}K^{bad})\right]~{}~{}~{}.$ (C.22)
As an aside, we simplify the triple derivative as follows
¯${\bar{\nabla}}_{a}{\bar{\nabla}}_{b}{\bar{\nabla}}_{c}K^{(ab)c}=$${\bar{\nabla}}_{a}{\bar{\nabla}}_{b}{\bar{\nabla}}_{c}K^{abc}+{\bar{\nabla}}_{b}{\bar{\nabla}}_{a}{\bar{\nabla}}_{c}K^{abc}$ (C.23)
$=$ $2{\bar{\nabla}}_{a}{\bar{\nabla}}_{b}{\bar{\nabla}}_{c}K^{abc}+{\bar{R}}^{a}{}_{dba}{\bar{\nabla}}_{c}K^{dbc}+{\bar{R}}^{b}{}_{dba}{\bar{\nabla}}_{c}K^{adc}$ (C.24)
where in the second line we have used the Bianchi identity for $\bar{R}^{a}{}_{bcd}$. The second and third terms cancel owing to the antisymmetry and symmetry properties ${\bar{R}}^{a}{}_{dba}=-{\bar{R}}^{a}{}_{dab}=-{\bar{R}}_{db}=-{\bar{R}}_{bd}$ and a relabeling of indices. Using the antisymmetry of $K^{abc}$ on the remaining first term, we again use the Bianchi identity to reduce this to
¯${\bar{\nabla}}_{a}{\bar{\nabla}}_{b}{\bar{\nabla}}_{c}K^{(ab)c}=$${\bar{\nabla}}_{a}({\bar{R}}^{a}{}_{dbc}K^{dbc}+{\bar{R}}^{b}{}_{dbc}K^{adc}+{\bar{R}}^{c}{}_{dbc}K^{abd})$ (C.25)
$=$ ¯${\bar{\nabla}}_{a}({\bar{R}}^{a}{}_{dbc}K^{dbc}+{\bar{R}}_{dc}K^{adc}-{\bar{R}}_{db}K^{abd})$ (C.26)
$=$ ¯${\bar{\nabla}}_{a}({\bar{R}}^{a}{}_{dbc}K^{dbc})$ (C.27)
where the second and third terms have vanishing owing to the symmetric $R_{dc}$ contracted over the antisymmetric indices of $K^{adc}$. Relabeling indices and inserting into Eq. (C.22) we have
¯${\bar{\nabla}}_{a}V^{a}=$ $\tfrac{1}{{\rm d}+1}{\bar{\nabla}}_{a}{\bar{\nabla}}_{b}\mathbb{D}^{ab}-\tfrac{\lambda_{0}^{2}}{{\rm d}+1}{\bar{\nabla}}_{a}\left({\bar{R}}^{a}{}_{bcd}K^{bcd}+2{\bar{{\mathcal{P}}}}_{db}K^{bad}\right)~{}~{}~{}.$ (C.28)
Using the definition of ${\mathcal{K}}^{a}{}_{bcd}$, we find the terms in the parenthesis reduce to ${\mathcal{K}}^{a}{}_{bcd}K^{bcd}$ thus we have simplified this to
¯${\bar{\nabla}}_{a}V^{a}=\tfrac{1}{{\rm d}+1}{\bar{\nabla}}_{a}{\bar{\nabla}}_{b}\mathbb{D}^{ab}-\tfrac{\lambda_{0}^{2}}{{\rm d}+1}{\bar{\nabla}}_{a}\left({\mathcal{K}}^{a}{}_{bcd}K^{bcd}\right)~{}~{}~{}.$ (3.70)